### v1

# Entangled photon pair excitation and time-frequency filtered multidimensional photon correlation spectroscopy as a probe for dissipative exciton kinetics

Abstract
In molecular aggregates, multiple delocalized exciton states interact with phonons, making the state-resolved spectroscopic monitoring of dynamics challenging. We propose a protocol that combines photon-entanglement-enhanced narrowband excitation of two-exciton states with time-frequency-filtered two-photon coincidence counting. It can alleviate bottlenecks associated with probing exciton dynamics spread across multiple spectral and temporal windows. We demonstrate that non-classical correlations of entangled photon pairs can be used to prepare narrowband two-exciton population distributions, circumventing transport in mediating states. The distributions thus created can be monitored using time-frequency-filtered photon coincidence counting, and the pathways contributing to photon emission events can be classified by tuning filtering parameters. Numerical simulations for a light-harvesting aggregate highlight the ability of this protocol to achieve selectivity by suppressing or amplifying specific pathways. Combining entangled photonic sources and multidimensional photon counting allow promising applications to spectroscopy and sensing.
I Introduction
In photosynthetic systems, light-harvesting complexes are critical components that facilitate efficient photon energy conversion through exciton formation and transport. Investigating the structural and functional mechanisms of exciton dynamics provides essential insights for engineering biomimetic energy-harvesting systems. In these systems, aggregated chromophores serve as the energy harvesters. The aggregation of individual chromophores results in delocalized electronic excitations: molecular excitons which are primarily of the Frenkel type. The excitons, often generated through photo-excitation, participate in multiscale relaxation and dephasing kinetics, due to the interaction with phonons [16, 1, 30], ultimately leading to the primary charge separation processes. Previous investigations have sought to monitor the role of coherent energy transfer that results from transition dipole-dipole interactions. The later can be probed via multi-pulse four-wave mixing techniques, popularly known as two-dimensional spectroscopy. The investigation multi-exciton states, where signatures of exciton-exciton correlations can be found, have comparatively received less attention. The early attributes of such correlations can already be seen in the two-exciton state energetics. Prior investigations have demonstrated that various modalities of two-photon absorption spectroscopy, resonant and off-resonant, can spectroscopically monitor these resonances [50, 51, 20, 15, 35, 54, 49, 7, 53, 26, 25, 33, 48, 21, 45, 29, 4]. Probing state-resolved dynamics in the two-exciton manifolds is complicated by the high density of states that take part in phonon-induced dissipation spanning multiple time and energy scales. Over the past several years, a series of proposals, limited to atomic, and simple molecular systems have underscored the advantages of employing entangled photon sources to improve accessibility of doubly excited states while scaling favorably with the probe intensity [46, 23, 44, 40, 14, 22, 8, 18, 19, 17, 43, 52, 32, 5, 34]. These proposals were complemented by examinations of the role of photon entanglement in the context of the reported enhancements [42, 28, 27, 39, 41, 24, 3, 6]. Parallel theoretical investigations have established that exploiting the non-classical time-frequency correlations of photonic states may yield superior signal resolution in spectroscopic measurements compared to classical counterparts [10, 8].
To monitor the fate of two-exciton states, multiphoton fluorescence intensity correlation measurements have been proposed [34, 31, 38, 11, 13]. In a series of work [11, 13], the concept of multidimensional signal measurements was extended to include coincidence counting measurements. This theoretical framework promises to transparently link information derived from photon correlation measurements to molecular events. The time and frequency filtered $n$-photon coincidence counting measurements provide a $2n$-dimensional parameter space for the events to be classified. However to firmly establish their role as a routine spectroscopic tool in extended molecular systems, it requires a protocol combining them with efficient initial state preparation strategies.
The two-exciton states in molecular aggregates, such as light-harvesting complexes, are typically prepared via the manifold of one-exciton states, which are subject to phonon-induced transport and dephasing. Furthermore, as the number of chromophoric sites increases, the number of one-exciton and two-exciton states increases as well. As a consequence, the number of interfering optical transitions may increase dramatically. Additionally, phonon-induced transport redistributes the population among the participating states during the dynamics.
In this work, we demonstrate that state-selective entangled photon excitation and time-frequency-filtered multidimensional correlation spectroscopy can serve as an efficient probe for dissipative exciton kinetics. Specifically, this approach helps unravel two-exciton transport kinetics by monitoring cascaded photon emission events, tracking the correlation between these events, and benchmarking the nature of the exciton wavepacket created by entangled photon excitation.
We focus on the light-harvesting complex LHCII, the most abundant antenna complex in plants and an important constituent of Photosystem II. It exists as a trimeric structure composed of structurally rigid monomeric aggregates; each monomer contains of several chlorophylls whose lowest energy levels are spectroscopically accessible. Exciton dynamics in LHCII exhibit significant complexity, particularly regarding the behavior of higher-order excitations. The combined protocol is shown to provide mechanistic insight into the cooperative exciton kinetics of LHCII, even in the presence of spectral weight redistribution caused by mode mixing and relaxation.
Figure 1: (A) Illustration of the proposed protocol, consisting of the driving, evolution, and detection stages. The process is assisted by entangled photon pairs and temporal-spectral filtering of emitted photons. In (B), three plausible snapshots of the dynamics (I)-(III), featuring increasingly complex pathways, that may be monitored using the protocol is displayed. In the simulation, the exciton transport within both manifolds is accounted for. Since these pathways may be simultaneously present in the dynamics, the protocol allows for selective probing and aids in constructing a mechanistic narrative.
In the following, in section II, we introduce the exciton-phonon Hamiltonian of the light-harvesting complex LHCII, which serves as our test system, and the photonic sources. This is complemented by the Appendix A. In the subsequent three subsections of Sec. III (III.1III.3), we detail the theoretical framework and numerical simulations of the combined protocol. The manuscript concludes in Sec. IV with a summary of our findings, a discussion of future outlook, and a brief examination of the current limitations.
II Driven dissipative one- and two-exciton kinetics
The dissipative one- and two-exciton dynamics of the light-harvesting complex LHCII are simulated using the Frenkel exciton model within the Heitler-London approximation. The model is constructed by considering the lowest three electronic energy states of each of the chlorophyll molecules. The Hamiltonian (with $\hbar=1$) is given by,
$H$ $=\sum_{mn}(E_{m}\delta_{mn}+J_{mn})B_{m}^{\dagger}B_{n}$
$+\sum_{mn}U_{m}^{(1)}B_{m}^{\dagger}B_{m}^{\dagger}B_{m}B_{m}+U_{mn}^{(2)}B_{m}^{\dagger}B_{n}^{\dagger}B_{n}B_{m}$
$+\sum_{j}\omega_{j}(b^{\dagger}_{j}b_{j}+\frac{1}{2})+\sum_{m,j}g_{m,j}(b^{\dagger}_{j}+b_{j})B_{m}^{{\dagger}}B_{m}$ (1)
Here, the on-site excitation energy, $E_{m}$, and inter-site hopping, $J_{mn}$, are accounted for by the first term in the Hamiltonian. The Coulomb interaction, responsible for the inter-site hopping, results in the delocalization of the exciton states across the sites. The higher-order excitonic nonlinearities that shift exciton energy states are given by $U_{m}^{(1)}$ and $U_{mn}^{(2)}$. The exciton creation (annihilation) operators, $B_{m}^{\dagger}$ ($B_{m}$), follow the commutation relation $B_{n},B_{m}^{\dagger}]=\delta_{mn}(1-\eta(B_{m}^{\dagger}B_{m})^{2})$, where $\eta=3/2$, while the phonon creation (annihilation) operators, $b_{j}^{\dagger}$ ($b_{j}$), follow $b_{i},b_{j}^{\dagger}]=\delta_{ij}$. The two-exciton Hamiltonian can be constructed as $H_{mnkl}^{(2)}=H_{mn}^{(1)}\delta_{kl}+\delta_{mn}H_{kl}^{(1)}$, where $H_{mn}^{(1)}$ is given by the first line of the Eq. II. For the LHCII, the number of chromophoric sites is $N_{s}=14$, which yields $N_{g}=1$ ground state, $N_{e}=14$ one-exciton states, and $N_{f}=105$ two-exciton states, composed of $14$ overtone and $91$ combination states. The excitonic eigenstates are obtained by the exact diagonalization of the exciton-number conserving Hamiltonian blocks as:
$|e_{j}\rangle$ $=\sum_{m}T^{(1)}_{j,m}B_{m}^{\dagger}|0\rangle$
$|f_{k}\rangle$ $=\sum_{mn}T^{(2)}_{k,mn}B_{m}^{\dagger}B_{n}^{\dagger}|0\rangle$ (2)
These transformation coefficients, $T^{(n)}$ (for the nthn-th manifold) define the spectral weights associated with the delocalized eigenstates. It is noteworthy that the energies of the two-exciton overtone and combination states, dependent on the nonlinearity terms $U_{m}^{(1)}$ and $U_{mn}^{(2)}$ respectively, deviate from the sum of two one-exciton states. As a consequence, the relative ordering of states within the two-exciton manifold is contingent upon two competing sets of parameters, governing exciton delocalization and localization. The latter influence, via excited state absorption processes, all three pathways responsible for the two-photon absorption events addressed in this article. The Green’s functions (valid for both one and two exciton manifold) describing can be represented in a compact form as follows
$\mathcal{G}_{a_{1}a_{2},a_{3}a_{4}}(t)=\delta_{a_{1}a_{2}}\delta_{a_{3}a_{4}}\theta(t)[\exp{(-Kt)}]_{a_{1}a_{1},a_{3}a_{3}}$
$+(1-\delta_{a_{1}a_{2}})\delta_{a_{1}a_{3}}\delta_{a_{2}a_{4}}\theta(t)\exp{(-i\omega_{a_{1}a_{3}}t-\gamma_{a_{1}a_{3}}t)}$ (3)
The optical properties of exciton aggregates consisting of multiple three level emitters are primarily dictated by the nature of intermolecular interactions and two-exciton nonlinearities [26, 2, 1]. The interaction with photonic sources, within the rotating wave approximation, can be described by the following Hamiltonian
$H_{\mathrm{int}}(t)=EV^{\dagger}+E^{\dagger}V=\sum_{j,pq\in\{ge,ef\}}\sqrt{\frac{2\pi\omega_{j}}{V}}$
$\Big(a_{j}\exp{(-i\omega_{j}t)}\exp{(i\omega_{pq}t)}B_{pq}+\mathrm{h.c.}\Big)$ (4)
where $B_{pq}(t)=d_{pq}|p\rangle\langle q|(\text{with properties}\,\,\{p,q\}\in g,e,f$, and $a_{j}$ are the exciton and photon annihilation operators, describing inter-manifold exciton transition and excitations in the photonic modes, respectively.
Figure 2: (A) Diagrammatic illustration (using the Schwinger-Keldysh loop representation) of the pathways involved in the proposed protocol, consisting of the driving, evolution, and emission stages. (B)-(C) Diagrammatic illustration (using the Liouville superoperator representation, suitable for dissipative systems) of the pathways involved in filtered emission (B) and excitation (C). The combinatorial pairing of five possible excitation diagrams and four possible emission diagrams gives rise to all possible pathways that contribute to the final signal. The corresponding key equations are presented in Section III.
For our case, the deployed photonic sources are entangled photon pairs generated via spontaneous parametric down-conversion (SPDC) in the weak-down-conversion limit. In a typical scenario, a birefringent crystal is pumped with a classical pump that precedes to the generation of correlated photon pairs with frequencies $\omega_{s},\omega_{i}$. It involves two field-matter interactions with the incoming classical field modes and four interactions with the outgoing signal-idler fields. The correlation properties of the emitted photon field is determined by the physical properties of the SPDC crystal [44, 47, 23, 46, 12, 8, 9]. The incoming entangled photon states can be described by the two photon state of the photon field
$\rho_{\text{photon}}$ $=\int\frac{d\omega_{1}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}^{\prime}}{2\pi}\int\frac{d\omega_{2}^{\prime}}{2\pi}$
$F(\omega_{1},\omega_{2})F^{*}(\omega_{1}^{\prime},\omega_{2}^{\prime})a_{\omega_{1}}^{{\dagger}}a_{\omega_{2}}^{{\dagger}}|0\rangle\langle 0|a_{\omega_{1}^{\prime}}a_{\omega_{2}^{\prime}}$ (5)
The entangled photon field correlation function is obtained as,
$\langle E^{{\dagger}}(\omega_{b}^{\prime})E^{{\dagger}}(\omega_{a}^{\prime})E(\omega_{b})E(\omega_{b})\rangle=F^{*}(\omega_{b}^{\prime},\omega_{a}^{\prime})F(\omega_{a},\omega_{b})$ (6)
where,
$F(\omega_{a},\omega_{b})$ $=\alpha A_{p}(\omega_{p})\Big\{\mathrm{sinc}[\phi(\omega_{a},\omega_{b})]$
$\exp{(i\phi(\omega_{a},\omega_{b}))}+\omega_{1}\leftrightarrow\omega_{2}\Big\}\text{with}$
$\phi(\omega_{a},\omega_{b})$ $=(\omega_{a}-\omega_{1})\tilde{T}_{1}/2+(\omega_{b}-\omega_{1})\tilde{T}_{2}/2$ (7)
where $\alpha$ governs the down-conversion efficiency. The temporal entanglement parameter of the photon pair sources is given by $\tilde{T}_{\mathrm{ent}}=\tilde{T}_{2}-\tilde{T}_{1}$ establish the upper bound for the interaction time interval. The transform-limited field profile of the SPDC pump, $A_{p}(\omega_{a},\omega_{b})$ is taken as a Gaussian, given by, $A_{p}(\omega_{a},\omega_{b})=E_{0}\sqrt{\pi/\Gamma_{p,0}}\exp{[-(\omega-\omega_{a}-\omega_{b})^{2}/4\Gamma_{p,0}]}$, where $\Gamma_{p,0}$ is the temporal width. The quantity is dependent on the time delays acquired by the signal and idler photons during the propagation within the nonlinear crystal. In contrast to entangled photon pairs, the correlation function of coherent laser fields would have factorized into the product of four amplitude functions, i.e.
$\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
$=A^{*}_{4}(\omega_{4})A^{*}_{3}(\omega_{3})A_{1}(\omega_{2})A_{1}(\omega_{1})$ (8)
and the signal would have scaled with the square of the incoming field intensity.
III Results
The excitation-evolution-emission events occur in three stages. Excitation by entangled photon pairs creates a tailored two-exciton wavepacket. The two-exciton wavepacket or density operator populations undergo transport, leading to a redistribution of populations over a finite waiting time. This redistributed population serves as the initial configuration for photon emission events. The cascaded two-photon emission process proceeds via two one-photon emissions interleaved by exciton transport in the one-exciton manifold. The signal can be expressed in a compact superoperator form
$S(\overline{t}_{s},\overline{\omega}_{s};\overline{t}_{i},\overline{\omega}_{i})$ $=\int_{-\infty}^{\infty}dt_{s}^{\prime}d\tau_{s}\int_{-\infty}^{\infty}dt_{i}^{\prime}d\tau_{i}\sum_{s,s^{\prime},i,i^{\prime}}\text{Tr}_{s,s^{\prime},i,i^{\prime}}[D^{(s,s^{\prime})}(\overline{t}_{s},\overline{\omega}_{s};t_{s}^{\prime},\tau_{s})D^{(i,i^{\prime})}(\overline{t}_{i},\overline{\omega}_{i};t_{i}^{\prime},\tau_{i})B^{(s,s^{\prime},i,i^{\prime})}(t_{s}^{\prime},\tau_{s};t_{i}^{\prime},\tau_{i})]$ (9)
where the superoperator for filtered detection
$D^{(\nu,\nu^{\prime})}(\overline{t}_{\nu},\overline{\omega}_{\nu};t_{\nu}^{\prime},\tau_{\nu})$
$=D(\overline{t}_{\nu},\overline{\omega}_{\nu};t_{\nu}^{\prime},\tau_{\nu})E_{\nu^{\prime}L}(t_{\nu}^{\prime})E_{\nu R}^{\dagger}(t_{\nu}^{\prime}+\tau_{\nu})$ (10)
and the bare two-photon signal
$B^{(s,s^{\prime},i,i^{\prime})}(t_{s}^{\prime},\tau_{s};t_{i}^{\prime},\tau_{i})$
$=T_{L}^{(s^{\prime},i^{\prime})}(t_{s}^{\prime},t_{i}^{\prime})T_{R}^{(s,i)\dagger}(t_{s}^{\prime}+\tau_{s},t_{i}^{\prime}+\tau_{i})$ (11)
is defined using the following superoperators corresponding to emission events
$T_{R}^{{(s^{\prime},i^{\prime})},{\dagger}}(t_{s}^{\prime},t_{i}^{\prime})$ $=i^{2}\int_{-\infty}^{t_{2}^{\prime}+\tau_{2}}dt_{4}\int_{-\infty}^{t_{1}^{\prime}+\tau_{1}}dt_{3}$
$E_{sR}(t_{3})E_{iR}(t_{4})V_{R}^{{\dagger}}(t_{4})V_{R}^{{\dagger}}(t_{3})$
$T_{L}^{(s^{\prime},i^{\prime})}(t_{s}^{\prime},t_{i}^{\prime})$ $=i^{2}\int_{-\infty}^{t_{2}^{\prime}}dt_{2}\int_{-\infty}^{t_{1}^{\prime}}dt_{1}$
$E_{sL}^{\dagger}(t_{1})E_{iL}^{\dagger}(t_{2})V_{L}(t_{1})V_{L}(t_{2})$ (12)
The averaging has been carried out over the two-exciton density operator following transport, i.e., $\rho_{f^{\prime}f^{\prime}}(t)=\mathcal{G}_{f^{\prime}f^{\prime},ff}(t-t^{\prime})\rho_{ff}(t^{\prime})$. Under a set of physically motivated approximations, the averaging can be carried out over each of these three terms independently. We also define the entangled photon driven excitation superoperator for future convenience, given by
$O(t)=T_{R}^{{\dagger}}(\tau_{2},\tau_{3})T_{L}(\tau_{1},\tau_{4})=i^{2}\int_{-\infty}^{\tau_{2}^{\prime}}d\tau_{4}\int_{-\infty}^{\tau_{1}^{\prime}}d\tau_{3}$
$\int_{-\infty}^{\tau_{3}}d\tau_{2}\int_{-\infty}^{\tau_{4}}d\tau_{1}E_{L}^{\dagger}(\tau_{4})E_{L}^{\dagger}(\tau_{1})E_{R}(t_{3})E_{R}(\tau_{2})$
$\qquad\qquad V_{R}^{{\dagger}}(\tau_{3})V_{R}^{{\dagger}}(\tau_{2})V_{L}(\tau_{4})V_{L}(\tau_{1})$ (13)
In the above, we assume that the excitation and emission events are not temporally overlapping, which results in the signal in Eq. 9 being factorized in terms of three principal contributions: first, preparation, involving exciton-photon (driving fields) interactions that lead to the preparation of the two-exciton density operator; second, evolution, describing two-exciton transport over a parametric time interval; and third, detection, describing exciton-photon (detection fields) interactions that yield the time-frequency filtered two-photon coincidence. In subsequent sections, we systematically describe all three contributions.
Figure 3: The two-exciton population distribution map, constructed by sequentially targeting each of the $105$ states in that manifold, using degenerate entangled photon pairs i.e., $\omega_{i}=\omega_{p}/2=E_{\mathrm{target}}^{(2)}/2$. The left panel serves as a reference simulation while the temporal width of the SPDC pump (entanglement time) is varied in the middle panel (right panel). Dispersal of the plot, in the horizontal cut, away from diagonal signifies a reduction in the narrowband excitation. For discussion, see section III.1
III.1 Entangled photon-pair excitation
The quantity of interest at the preparation stage, the two-exciton population distribution, can be defined using a projection operator acting on the diagonal matrix elements of the phonon-averaged exciton density operator, expressed as: $\rho_{ff}(t)=\text{Tr}[O(t)|f\rangle\langle f|]_{\rho_{0}}$. Here, a factorizable initial condition between the driving, the detection field modes, and the exciton density operator is assumed. We perturbatively expand the combined photon-exciton-phonon density operator to the fourth order in $H_{\text{int}}(t)$, and partially trace over the phonon modes to obtain the desired expression for the observable (derivation detailed in Appendix B) as follows:
$\rho_{ff}(t)$ $\propto 2\text{Re}\Big\{\int\frac{d\omega_{4}}{2\pi}\int\frac{d\omega_{3}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}}{2\pi}e^{+i\omega_{3}\tau_{3}+i\omega_{4}\tau_{4}-i\omega_{2}\tau_{2}-i\omega_{1}\tau_{1}}\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
$\Big\{d_{fe^{\prime}}d_{e^{\prime}g}d_{fe}^{*}d_{eg}^{*}\times I(-\omega_{4}-\omega_{3}+\omega_{2}+\omega_{1};-z_{ff})I(-\omega_{3}+\omega_{2}+\omega_{1};-z_{fe^{\prime}})I(+\omega_{2}+\omega_{1};-z_{fg})I(\omega_{1};-z_{eg})$
$+d_{fe^{\prime}}d_{fe}^{*}d_{e^{\prime}g}d_{eg}^{*}\times I(-\omega_{4}+\omega_{2}-\omega_{3}+\omega_{1};-z_{ff})I(\omega_{2}-\omega_{3}+\omega_{1};-z_{fe^{{}^{\prime\prime}}})$
$\left(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}I(-\omega_{3}+\omega_{1};-z_{p})\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}I(-\omega_{3}+\omega_{1};-z_{ee^{\prime}})\right)I(\omega_{1};-z_{eg})$
$+d_{fe^{\prime}}d_{fe}^{*}d_{e^{\prime}g}d_{eg}^{*}\times I(+\omega_{2}-\omega_{4}-\omega_{3}+\omega_{1};-z_{ff})I(-\omega_{4}-\omega_{3}+\omega_{1};-z_{fe^{\prime\prime}})$
$\left(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}I(-\omega_{3}+\omega_{1};-z_{p})\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}I(-\omega_{3}+\omega_{1};-z_{ee^{\prime}})\right)I(\omega_{1};-z_{eg})\Big\}\Big\}$ (14)
It is a sum of five distinct terms that represent interfering exciton pathways (see Liouville space Feynman diagrams presented in Fig. 2 for more insights). It includes contributions from both one- and two-exciton coherence and population terms of the exciton density operator. The first term represents the two-exciton coherence-mediated pathway, which is susceptible to inter-manifold dephasing. The second and fourth terms describe pathways susceptible to one-exciton population transport. The third and fifth terms describe pathways susceptible to intra-manifold dephasing. The convolutional form is particularly advantageous for introducing the field correlation functions, given in Eq. 6-Eq. II. The field correlation functions can be manipulated to selectively suppress and amplify specific exciton resonances. Next, we illustrate these capabilities via numerical simulations for both degenerate and nondegenerate entangled photon pairs.
III.1.1 Simulation: excitation of the full suite of two-exciton states using degenerate entangled photon pairs
In this section, we monitor the final population distribution following the excitation via spectrally degenerate entangled photon pairs, i.e., $\omega_{1}=\omega_{2}=\omega_{p}/2=E_{\mathrm{target}}^{(2)}/2$. We had carried out a parametric scan, taking each of the $N_{f}=105$ two-exciton states as a target, and present the final normalized population distribution in Fig. 3. The simulation presented in the left column uses an entanglement time of $T_{\mathrm{ent}}=10\,\mathrm{fs}$ and an SPDC pump width of $\tau_{0}=150\,\mathrm{fs}$, and acts as a reference case. Variations in the temporal width of the SPDC pump ($\tau_{0}=50\,\text{fs}$) and the entanglement time parameter ($T_{\text{ent}}=60\,\text{fs}$) are presented in the middle and the right column, respectively. We note that target states near the middle of the energy ladder exhibited broader distributions around the target in all three cases. In other words, we find a lower degree of selectivity for target states in the middle rung of the two-exciton manifold and higher selectivity for those belonging to the lower rung. Altering the width of SPDC pump pulse in the middle column resulted in diminished selectivity, likely due to the activation of multiple pathways. Altering the entanglement time in the right column preserved the selectivity for the majority of states. Across all plots, we notice a trend: some two-exciton states remain preferentially populated across all parameter regimes. This feature have likely emerged from an intricate interplay of exciton transport parameters leading to the absence of specific relaxation channels.
Figure 4: The population distribution for the target state $f_{07}$ which is dominantly excited via the intermediate states $e_{07}$ and $e_{09}$ is displayed for variations in the entangled photon parameters (rows) and for two distinct transport times (columns): $t=0\,\text{fs}$ (immediately after the pulse) and $t=50\,\text{fs}$. In the upper panel, the entanglement time is $\tilde{T}_{\text{ent}}=10\,\text{fs}$ and the SPDC pump width is $\tau_{0}=150\,\text{fs}$. In the middle panel (lower panel), the width is reduced (entanglement time is increased) to $\tau_{0}=50\,\text{fs}$ ($\tilde{T}_{\text{ent}}=30\,\text{fs}$). For detailed discussion, refer to Section III.1.
III.1.2 Simulations: differential preparation of two-exciton populations via preselected mediating states and parametric dependence on photonic properties
Now, we focus on two specific two-exciton states, $f_{07}$ and $f_{83}$, among the full suite of target states previously monitored, and examine the possibility of their excitation via a chosen pair one-exciton states, selected by setting $\omega_{1}=E_{07}$ and $\omega_{2}=E_{09}$. The mediating states are delocalized over sites and prone to ultrafast transport. The results are displayed in the left column of Fig. 4, and Fig. 5, respectively. The entanglement time and width of SPDC pump are set at $T_{\mathrm{ent}}=10\,\mathrm{fs}$ and $\tau_{0}=150\,\mathrm{fs}$, for the upper panel (similar to the reference case in Fig. 3). The width is varied to $\tau_{0}=50\,\mathrm{fs}$ in the middle panel, and the entanglement time is varied to $T_{\mathrm{ent}}=40\,\mathrm{fs}$ in the bottom panel. The results demonstrate that the desired narrowband excitation is indeed possible, albeit with low final occupancy. The increase in the temporal width results in a significant loss of selectivity as the number of mediating pathways with comparable spectral weights grows. The variation of the entanglement time didn’t result in any major changes, except for the tailoring of the relative populations. The latter, however, will have consequences as the population redistributes over time, as will be shown in the next section.
III.2 Transport in the two-exciton manifold
The created population distributions undergo time evolution governed by the parameters of the two-exciton transport matrix. Here, the following questions are worth investigating: What is the nature of the population evolution for a given state distribution? How do two initially created two-exciton population distributions undergo redistribution? And how does the nature of redistribution kinetics vary with driving field parameters?
The kinetic equation that governs the dissipative time evolution of exciton populations, i.e., exciton transport is given by
$\frac{d}{dt}\rho_{ff}(t)=-\sum_{f^{\prime}}K_{ff,f^{\prime}f^{\prime}}\rho_{f^{\prime}f^{\prime}}(t)$ (15)
where the transport rate matrices, $K_{ff,f^{\prime}f^{\prime}}$ follow the detailed balance condition $K_{f^{\prime}f^{\prime},ff}/K_{ff,f^{\prime}f^{\prime}}=\exp(-\omega_{f^{\prime}f}/(k_{B}T))$, and $\sum_{f}K_{ff,f^{\prime}f^{\prime}}=0$. The initial two-exciton population distribution is obtained via the Eq. III.1 which is propagated using the expression: $\rho_{ff}(t_{2})=G_{ff,f^{\prime}f^{\prime}}^{(N)}(t_{2}-t_{1})\rho_{f^{\prime}f^{\prime}}(t_{1})$, where the $G_{f^{\prime}f^{\prime},ff}^{(N)}(\tau)$ is given by first term in Eq. II. The following four sets of simulations will help us address the questions posed above.
III.2.1 Simulations: differential evolution of two-exciton populations and parametric dependence on photonic properties
We track the differential population evolution following the excitation targeted at $f_{07}$ (right column in Fig. 4) and $f_{83}$ (right column in Fig. 5), taken at a delay of $t=50\,\text{fs}$, and $t=250\,\text{fs}$, respectively. Each of the three panels, upper, middle and bottom, represents simulation using different photonic parameters. The parameters were listed in the preceding section III.1.2 where we discussed the reference simulation. We observe that the evolved distributions at short times retained the memory of prepared distribution in Fig. 4, but gets washed out in Fig. 5. The effect of the variation of photonic parameters is also less visible in the latter. The rapid redistribution and localization towards a few selected states in the latter is a key feature that requires future investigation. It can also be concluded that such a prepared distribution would loose the principal advantages of the entangled photon pair excitations.
Figure 5: The population distribution for the target state $f_{83}$ which is dominantly excited via the intermediate states $e_{07}$ and $e_{09}$ is displayed for variations in the entangled photon parameters (rows) and for two distinct transport times (columns): $t=0\,\text{fs}$ (immediately after the pulse) and $t=250\,\text{fs}$. In the upper panel, the entanglement time is $\tilde{T}_{\text{ent}}=10\,\text{fs}$ and the SPDC pump width is $\tau_{0}=150\,\text{fs}$. In the middle panel (lower panel), the width is reduced (entanglement time is increased) to $\tau_{0}=50\,\text{fs}$ ($\tilde{T}_{\text{ent}}=30\,\text{fs}$). For detailed discussion, refer to Section III.1.
III.2.2 Simulations: differential redistribution of two-exciton populations
In Fig. 6, we track the population redistribution by displaying the population snapshots at $t=50,100,250,1000\,\text{fs}$ following the excitation targeted at $f_{07}$ (in the left column) and $f_{83}$ (in the right column). The results demonstrate two distinct features: the population migrates to the lower energy sector of the manifold rather monotonically for $f_{07}$ and by a slower intra-manifold cascading mechanism for $f_{83}$. The slower transport can be attributed to lack of overlap between the latter with the state in the middle rung of two-exciton manifold.
The population redistribution can be spectroscopically monitored by combining spectral-temporal filtering schemes in the detection of the two-photon emission. [12, 10, 36]. Such a scheme will track cascaded inter-manifold transitions fluorescence that competes with the intra-manifold transport. The next section is devoted to the theory and simulations demonstrates these schemes.
Figure 6: Population redistribution at four different time points following the excitation of the target states $f_{07}$ (right column) and $f_{83}$ (left column). The results demonstrate key features of the dynamics, including rapid funneling of the population along energy gradients and cascaded population migration for the two targets, respectively. For detailed discussion, refer to Section III.2
III.3 Multipoint photon correlation spectroscopy
The quantity of interest in photon correlation spectroscopy is the time-frequency filtered two-photon coincidence counting signal. In this section, we describe how this quantity can monitor the time-evolved population distributions that serve as the initial configuration for the photon emission pathways. The signal can be expressed as:
$S(\{\bar{t}_{j},\bar{\omega}_{j}\})=\int dt_{1}^{\prime}\int d\tau_{1}D^{(1)}(t_{1},\omega_{1};t_{1}^{\prime},\tau_{1})$
$\times\int dt_{2}^{\prime}\int d\tau_{2}D^{(2)}(t_{2},\omega_{2};t_{2}^{\prime},\tau_{2})\langle N_{r}(t_{1}^{\prime},\tau_{1})N_{s}(t_{2}^{\prime},\tau_{2}^{\prime})\rangle$
$=\int dt^{\prime}_{1}\int d\tau_{1}\int dt^{\prime}_{2}\int d\tau_{2}D^{(1)}(t_{1}^{\prime},\tau_{1})D^{(2)}(t_{2}^{\prime},\tau_{2})$
$\sum_{s,s^{\prime}}\sum_{i,i^{\prime}}\langle E_{s^{\prime}}^{\dagger}(t^{\prime}_{2}+\tau_{2})E_{i^{\prime}}^{\dagger}(t^{\prime}_{1}+\tau_{1})E_{i}(t^{\prime}_{1})E_{s}(t^{\prime}_{2})\rangle$ (16)
where $N(t^{\prime},\tau)=\sum_{s,s^{\prime}}\hat{E}_{sR}^{\dagger}(t^{\prime}+\tau)\hat{E}_{s^{\prime}L}(t^{\prime})$ is a bare photon number operator and $E_{s}(t)=\sqrt{2\pi\hbar\omega_{s}/\Omega}\hat{a}_{s}e^{-i\omega_{s}t}$, and $\Omega$ is the mode quantization volume corresponding to the detector modes. We also defined two detector spectrogram functions: $D^{(j)}(\bar{t}_{j},\bar{\omega}_{j};t_{j}^{\prime},\tau_{j})$ that are governed by the parameters $\bar{t}_{j}$ and $\bar{\omega}_{j}$, describing the detector centering times for the temporal, and spectral filter, respectively. It can be expressed in terms of the filtering functions as,
$D(\bar{t},\bar{\omega},t^{\prime},\tau)$
$=\int\frac{d\omega^{\prime\prime}}{2\pi}e^{-i\omega^{\prime\prime}\tau}|F_{\bar{\omega}}(\omega^{\prime\prime},\omega)|^{2}F_{\bar{t}}^{*}(t^{\prime}+\tau,t)F_{\bar{t}}(t^{\prime},t).$ (17)
Modulation of these functions individually allows for the examination of the photon emission pathways and the systematic recovery of the photon correlation functions. For this work, we choose the Lorentzian filtering functions (see Appendix D), which leads to the spectrogram being expressed as:
$D(\bar{t},\bar{\omega},t^{\prime},\tau)=\frac{1}{2\sigma_{\omega}}\theta(t^{\prime}-t)\theta(t^{\prime}+\tau-t)$
$[\theta(\tau)e^{-\sigma_{\bar{\omega}}\tau}+\theta(-\tau)e^{\sigma_{\bar{\omega}}\tau}]e^{-(i\omega+\sigma_{\bar{t}})\tau-2\sigma_{\bar{t}}(t^{\prime}-t)}$ (18)
We lay out the filtering strategy: the spectral filters allow state-resolution for the dynamics occurring within similar transport windows; the temporal filters help separate the transport processes that involve transitions with similar energy gaps; combined together, they allow for the tracking of individual state transitions. The filtering properties are controlled by the width of the temporal (spectral) filters, $\sigma_{\bar{\omega}_{j}}(\sigma_{\bar{t}_{j}})$, centering times of the temporal (spectral) gate, $\bar{t}_{j}$ ($\bar{\omega}_{j}$). The final signal can be expressed as the sum of four pathways (which is reduced to two following symmetry arguments):
$S(\{\bar{t}_{j},\bar{\omega}_{j}\})$ $=\sum_{p=1}^{4}S^{(p)}(\{\bar{t}_{j},\bar{\omega}_{j}\})=2\text{Re}\sum_{p=1}^{2}S^{(p)}(\{t_{j},\omega_{j}\})$
$=\sum_{g,e,e^{\prime},f,f^{\prime}}|\mathcal{D}(\omega_{1})\mathcal{D}(\omega_{2})|^{2}|d_{eg}|^{2}|d_{ef}|^{2}\times\int dt_{1}^{\prime}\int d\tau_{1}D_{>}^{(1)}(\bar{t}_{1},\bar{\omega}_{1};t_{1}^{\prime},\tau_{1})\times\int dt_{2}^{\prime}\int d\tau_{2}D_{>}^{(2)}(\bar{t}_{2},\bar{\omega}_{2};t_{2}^{\prime},\tau_{2})$
$\langle\langle I|\mathcal{G}_{ge^{\prime},ge^{\prime}}(\tau_{1})\mathcal{G}_{e^{\prime}e^{\prime},ee}(t_{1}^{\prime}-t_{2}^{\prime}-\tau_{2})\mathcal{G}_{ef^{\prime},ef^{\prime}}(\tau_{2})\mathcal{G}_{f^{\prime}f^{\prime},ff}(t_{2}^{\prime})|\rho_{ff}\rangle\rangle$
$+\sum_{g,e,e^{\prime},f,f^{\prime}}|\mathcal{D}(\omega_{1})\mathcal{D}(\omega_{2})|^{2}|d_{eg}|^{2}|d_{ef}|^{2}\times\int dt_{1}^{\prime}\int d\tau_{1}D_{>}^{(1)}(\bar{t}_{1},\bar{\omega}_{1};t_{1}^{\prime},\tau_{1})\times\int dt_{2}^{\prime}\int d\tau_{2}D_{<}^{(2)}(\bar{t}_{2},\bar{\omega}_{2};t_{2}^{\prime},\tau_{2})$
$\langle\langle I|\mathcal{G}_{ge^{\prime},ge^{\prime}}(\tau_{1})\mathcal{G}_{e^{\prime}e^{\prime},ee}(t_{1}^{\prime}-t_{2}^{\prime})\mathcal{G}_{f^{\prime}e,f^{\prime}e}(-\tau_{2})\mathcal{G}_{f^{\prime}f^{\prime},ff}(t_{2}^{\prime}+\tau_{2})|\rho_{ff}\rangle\rangle$ (19)
The terms $D_{>}$ and $D_{<}$ indicate the sign of the delay $\tau$ variables, and $\mathcal{D}(\omega_{1})\mathcal{D}(\omega_{2})$ is the density of states for the detector, taken at an idealized, narrowband limit around each frequency $\omega_{j}$. In our simulation, the influence of this delay was accounted for by evaluating the action of the transport Green’s functions on the initial states within the respective manifolds.
The rationale for deploying such a scheme can be developed in the following manner. During population transport, the ensuing redistribution erases the signatures of the initially created two-exciton wavepacket. The timescales of this phenomena depend on the properties of the transport Green’s functions. The redistribution give rise to new inter-manifold dipole-allowed transitions and render some transitions diminished in spectral intensity.
Figure 7: The time-frequency filtered two-photon coincidence counting signal is displayed, showing variations in the widths of both filtering functions and the centering time of the temporal filtering function. To control the signal resolution, the following systematic variations relative to the reference simulation (upper-left) are presented: the width of the spectral filter (upper-right); the centering time of the temporal filtering function (middle-left); both the spectral filter width and the temporal filter centering time (middle-right); the width of the temporal filter (bottom-left); and the centering time of the temporal filtering function (bottom-right). The targeted excitation of $f_{07}$ by the entangled photon pairs serves as the initial configuration, which is followed by a subsequent series of events: two-exciton manifold transport, inter-manifold transitions, one-exciton manifold transport, and subsequent inter-manifold transitions. For a diagrammatic illustration of the processes involved, see Fig. 2, and for detailed discussion, refer to Section III.3.1
The subsequent photon emission events can be monitored by appropriately selecting the spectral and temporal filtering properties of the first set of detectors. Following the photon emission, the dynamics continue in the one-exciton manifold where transport population scrambling takes place for a time interval, followed by another photon emission event to the ground state. Three representative pathways in Fig.1 shows complexities that may be present.
The joint spectral-temporal filtering strategy can be a powerful tool that enhances the resolution of the bare photon coincidence counting signal. Two sets of spectral-temporal filters, defined by the spectrogram can accomplish two distinct tasks: first, the photon emissions events can be temporally windowed to focus on transport timescales and spectrally windowed to focus on the relevant exciton states; second, the photon emissions events can be spectrally windowed such that pairs of inter-manifold transitions originating from the similar two-exciton transport events can be monitored. Monitoring these events allows us to construct a kinetic description of dissipative exciton dynamics events triggered by the entangled photon excitation. It also help infer how long the remnants of the initially created wavepacket persist. The diagrams presented in Figure 2 clarify the physical processes in a more transparent manner.
III.3.1 Simulation: two-photon coincidence counting signal with variation in filtering parameter and different waiting times
In Fig. 7, we present six two-dimensional plots illustrating the signal where the filtering widths and centering parameters (the latter relates to the transport times) are varied. The centering times of the spectral filters are taken as the scanning parameters. In these plots, a series of features can be noticed: the peaks along the $\bar{\omega}_{1}$ axis correspond to transitions from the two to and one-exciton manifold ($\sim\omega_{fe}$), while the peaks along the $\bar{\omega}_{2}$ axis correspond to transitions from the one-exciton manifold to the ground state ($\sim\omega_{eg}$). The one-exciton manifold has a spectral width of $\approx 950\,\text{cm}^{-1}$. The density of energy gaps is significantly higher in the upper set of transitions. Even though, for the chosen example, the initial state distribution, following excitation of $f_{07}$, effectively spans only the first ten levels of the two-exciton manifold (an energy range of $\approx 350\,\text{cm}^{-1}$), the number of possible $\omega_{fe}$ transitions remains high.
The upper-left panel (the reference simulation for this set) employs the following filtering parameters: spectral widths $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=10\,\text{cm}^{-1}$, temporal widths $\sigma_{\bar{t}_{1}}=\sigma_{\bar{t}_{2}}=4.8681\,\text{cm}^{-1}$, and two-exciton (one-exciton) waiting times $t_{\text{w},1}=0\,\text{fs}$ ($t_{\text{w},2}=100\,\text{fs}$). Variables $t_{\text{w},j}$ are associated with the $t_{2},t_{1}$ integrals when the detector spectrograms are inserted into the relevant expression. We observe eight resonance zones spread along the $\bar{\omega}_{1}(/10^{3}\,\text{cm}^{-1})$ axis, each containing several resonances that can be resolved along the $\bar{\omega}_{2}(/10^{3}\,\text{cm}^{-1})$ axis. For easier interpretation, we may classify these resonances along $\bar{\omega}_{1}$ as four distinct clusters, containing the following number of resonance zones: two ($\sim 14.0-14.2$), three ($\sim 14.3-14.7$), two ($\sim 14.9-15.1$), and one ($\sim 15.3-15.5$). Each of these clusters may have multiple origins in the upper manifold. The four observed $\bar{\omega}_{1}$ resonance clusters correspond approximately to the transitions $\omega_{f_{1}e_{13}}$ and $\omega_{f_{2}e_{10}}-\omega_{f_{2}e_{13}}$; $\omega_{f_{1}e_{8}},\omega_{f_{2}e_{6}},\omega_{f_{2}e_{5}}$; $\omega_{f_{4}e_{3}}$ and either $\omega_{f_{3}e_{1}}$ or $\omega_{f_{4}e_{1}}$; and $\omega_{f_{7}e_{1}}-\omega_{f_{8}e_{1}}$, respectively. Taking each cluster, the $\bar{\omega}_{2}$ ($10^{3}\,\text{cm}^{-1}$) resonances are located at $\sim 15.5,15.8$ and $\sim 15.4,15.7$ (corresponding to $\omega_{e_{8}g},\omega_{e_{12}g}$ and $\omega_{e_{8}g},\omega_{e_{12}g}$); $\sim 15.3,15.6$, $\sim 15.3,15.4$, and $\sim 15.3,15.4$ ($\omega_{e_{7}g},\omega_{e_{8}g}$; $\omega_{e_{7}g},\omega_{e_{8}g}$; and $\omega_{e_{7}g},\omega_{e_{8}g}$); $\sim 15.1,15.3$ ($\omega_{e_{1}g},\omega_{e_{3}g}$); and $\sim 15.0$ ($\omega_{e_{1}g}$), respectively. This plot will serve as a reference simulation.
The upper-right panel presents the results of varying the spectral filter width by setting $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$, while keeping all other parameters fixed to those of the reference simulation. We observe that several resonance zones in the second cluster are absent along the $\bar{\omega}_{1}$ axis, with the most noticeable absences occurring in the middle and lower energy sectors. Additionally, among the present zones, a few resonances along the $\bar{\omega}_{2}$ axis are absent, and the fourth cluster has been diminished in intensity. This observation may be rationalized by noting the increased destructive interferences permitted by the broadband spectral filter.
The middle-left panel investigates the effect of longer waiting times, specifically using $t_{\text{w},2}=1000\,\text{fs}$ for the one-exciton transport, while keeping all other parameters unchanged from the reference simulation. The population distribution in the one-exciton manifold is dependent on the cascaded emission from the two-exciton manifold. We note that the clusters have further coalesced along the $\bar{\omega}_{1}$ axis and are clearly separated into two dominant regions. The latter is indicative of the dipole transition rule dominating the spectral weight at these longer times. Furthermore, the fourth cluster has vanished, and the resolution along the $\bar{\omega}_{2}$ axis is lost. The middle-right panel explores the variation of resolution compared to the middle-left one by varying both the spectral filter width and the one-exciton waiting time ($t_{\text{w},2}$). The parameters are set jointly equal to those used in the two previous simulations ($\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$ and $t_{\text{w},2}=1000\,\text{fs}$), while keeping all other parameters fixed. We observe that widening the spectral gate may not have permitted a larger number of resonance structures along the $\bar{\omega}_{2}$ axis. This implies that a broader set of transitions remain unaccounted for (either not presented or not dipole allowed). Importantly, these transitions, despite differing widely in energy, participate in similar dynamics due to the longer transport window afforded by the extended waiting time.
The lower panels are dedicated to exploring two distinct cases: the bottom-left panel examines the role of the width of the temporal filter (by setting $\sigma_{\bar{t}_{2}}=0.5409\,\text{cm}^{-1}$ ), and the bottom-right panel examines the role of the two-exciton transport waiting time (by setting $t_{\text{w},1}=50\,\text{fs}$), while keeping other parameters the same as the reference. For the bottom-left panel, we note that the second cluster has a diminished presence, with several resonances along the $\bar{\omega}_{2}$ axis being absent, an effect attributed to the shorter temporal filter effectively skipping some transitions occurring at longer respective timescales. For the bottom-right panel, in contrast with the reference simulation, we observe that the double resonance structures along the $\bar{\omega}_{2}$ axis have coalesced into single structures, while there is no significant change along the $\bar{\omega}_{1}$ axis. This can be attributed to the fact that at longer two-exciton transport duration, state-specific transport rates exhibited minor differences. The redistributed the spectral weights, in turn, has resulted in a fewer $\omega_{fe}$ transitions becoming dominant.
The initial states in all cases can be tracked from Figure 6. However, the state indices require a conversion to the energy domain (energy gaps) for constructing a robust phenomenological description.
IV Conclusions
In this article, we had demonstrated that a combined protocol utilizing entangled photon-pair excitation and temporally-spectrally filtered photon counting provides a systematic route to probe dissipative multi-exciton kinetics. The protocol, illustrated via simulations for the exciton aggregate LHCII, is presented across three subsequent sections. It revealed, in section III.1, that narrow-band distribution around a special class of two exciton states can indeed be excited with considerable degree of control. The final population distribution is benchmarked against the variation of the entangled photon parameters in section III.1.2. Here we examined the pathway selectivity by targeting two different two exciton state that are excited via two pre-selected one-exciton states. The exciton-phonon interactions in multi-site molecular aggregates are characterized by exciton transport and dephasing dynamics spanning multiple time and energy scales. The entangled photon pairs demonstrated a superior ability maintain pathway selectivity. Following the excitation, and population evolution, illustrated in section III.2.1-III.2.2), we introduced the detection scheme to monitor the fate of the time-evolving population, and classify the emission pathways. It was shown, in section III.3, that the joint spectral and temporal filtering of the emitted fluorescence allow a more nuanced monitoring of inter-manifold transitions competing with intra-manifold transport kinetics.
We note that the targeted states, in III.1.2, are chosen due to their higher overlap with the local overtone states. These properties can be quantified by analyzing the transformation matrix elements, $T^{(2)}_{k,mm}$. While state-selective excitations in interacting molecular aggregates conventionally yield spatially delocalized wavepackets, an important corollary suggests that a specially crafted exciton wavepacket, that is, a tailored superposition of exciton eigenstates, may allow for the excitation of spatially localized wavepackets. Such localization could then enhance the propensity of exciton-exciton annihilation provided that significant non-adiabatic coupling elements between the higher-lying states allow the internal conversion to occur.
The expression in Eq. III.1 remains valid for a wide range of correlated optical fields, albeit in the weak-field regime. In the expression, five terms appear, each of which contains a convolution of inter-manifold coherence and intra-manifold population Green’s functions with four-point photon correlation functions. The latter can be interpreted as pathway weights, and multiobjective optimization schemes can be used to actively design them. Future work will address this issue.
Similarly, the expressions in Eq. III.3 remain valid for any general initial state distributions. The variation of the filtering parameters offers us the ability to classify inter-manifold transitions in the presence and competing with intra-manifold transport kinetics. It allows better assignment timescales and generate a dynamical picture for one and two-exciton transport kinetics. We assumed that the excitation and emission events are temporally separated, omitting spontaneous Raman contributions.
From these simulations, a probing strategy may emerge: preferentially accessing the one-exciton resonances that are less susceptible to transport within a given temporal window may be the key to achieving controlled excitation. We note that the full protocol encountered transport events thrice during its course. Within the model and the chosen expressions, the entanglement time, controlled by the length of the birefringent crystal, determines the temporal window of the photon-exciton interactions. The bandwidth of the classical SPDC pump, on the other hand, determines the range of accessible resonances. The independent controllability of these variables affords some flexibility in selecting their values. In turn, it will allow control over a broad range of time and frequency parameter domain. Analogously, the choice of filtering parameters, decided independently, must take into account the excitation properties to be able to characterize a wide range of two-photon emission pathways.
The utility of correlated photonic sources for multiphoton spectroscopies promises two main advantages in the appropriate parameter regime: the ability to actively modulate multiple material pathways to highlight unseen spectral signatures, and favorable scaling of the signal with the incoming field intensity. However, these benefits face a significant challenge due to the inherently weak signal-to-noise ratio. Technological advances in this direction would make the proposed protocol a suitable investigative tool for exploring potential non-classical effects in light-harvesting complex. Alternatively, the scheme remains applicable in the high-gain regime using squeezed photon sources.
Acknowledgements.
A.D. acknowledges the support from DESY (Hamburg, Germany), a member of the Helmholtz Association HGF. S.M. gratefully acknowledges the support of the National Science Foundation (NSF) grant Grant No. CHE-2246379.
Appendix A The Hamiltonian
In the Hamiltonian, the Eq. II, the vibrational terms are specified in the third line. The normal modes associated with collective intermolecular vibrations constitute the phonon modes. The latter, in the continuum limit, i.e., $J_{\mathrm{ph}}(\omega)=\sum_{j}|g_{j}|^{2}[\delta(\omega-\omega_{j})-\delta(\omega+\omega_{j})]$ are described using an overdamped oscillator and $48$ multimode Brownian oscillators. The corresponding spectral function is given by
$J_{\mathrm{ph}}(\omega)$ $=2\lambda_{0}\frac{\omega\gamma_{0}}{\omega^{2}+\gamma_{0}^{2}}+\frac{2\lambda_{j}\omega_{j}^{2}\omega\gamma_{j}}{(\omega_{j}^{2}-\omega^{2})^{2}+\omega^{2}\gamma_{j}^{2}}$ (20)
Here, $\lambda_{0}$ and $\gamma_{0}$ represent the damping strength and relaxation parameter for the overdamped mode, respectively. The Brownian oscillators are characterized by the Huang-Rhys parameter, mode frequency, and correlation time, denoted as $\lambda_{j}$, $\gamma_{j}$, $\omega_{j}$, respectively. Relevant parameter values are detailed in the [9, 8].
The phonon interaction gives rise to exciton transport (first term) and dephasing (second term) Eq. II. The exciton transport components can be expressed as
$[\exp{(-K^{(n)}t)}]_{a_{1}a_{1},a_{3}a_{3}}$ $=\sum_{p}\chi_{a_{1}p}^{R}D_{pp}^{-1}e^{-\lambda_{p}t}\chi_{pa_{3}}^{L}$ (21)
where $\lambda_{p}$ is the p-th eigenvalue of the transport matrix, $\chi^{L}(\chi^{R})$ are the left (right) eigenvectors $K\chi^{R}=\lambda_{p}\chi^{R}$ and $\chi^{L}K=\lambda_{p}\chi^{L}$ of the same with additional relation $D_{pp}=\chi^{L}\chi^{R}$. The inter-manifold coherence green’s functions, contains excitonic resonances $i\omega_{a_{1}a_{2}}^{(n)}-\gamma_{a_{1}a_{2}}^{(n)}t=z_{a_{1}a_{2}}^{(n)}$, where the line-broadening function term is evaluated as
$\gamma_{a_{j}}^{(n)}(t)$ $=\sum_{a_{j^{\prime}}}C_{\mathrm{ph}}(\omega_{a_{j}a_{j^{\prime}}}^{(n)})\sum_{m}T_{ma_{j}}^{(n)}T_{ma_{j}}^{(n)}T_{ma_{j^{\prime}}}^{(n)}T_{ma_{j^{\prime}}}^{(n)}$ (22)
and the phonon correlation function evaluated at resonant exciton transition frequencies is given by
$C_{\text{phonon}}(\Omega)=\int_{0}^{\infty}dt\exp(i\Omega t)$
$\qquad\times\int\frac{d\omega}{2\pi}J(\omega)\Big[\coth{(\beta\omega/2)}\cos{\omega t}\mp i\sin{\omega t}\Big]$ (23)
Therefore characteristic broadening of the resonances is determined by the phonon spectral function, and exciton gaps.
Appendix B Note on the observable: two-exciton population distribution
The observable is defined using the projection operator onto the desired two-exciton state, averaged over the time-dependent density operator. Assuming the perturbative regime of exciton-driving field interaction to hold, the density operator can be expanded to the fourth-order in incoming field-exciton interaction Hamiltonian to obtain
$O_{ff}(t)$ $=(-i)^{4}\text{Re}\int^{\infty}_{-\infty}d\tau_{4}\int^{\tau_{4}}_{-\infty}d\tau_{3}\int^{\tau_{3}}_{-\infty}d\tau_{2}\int^{\tau_{4}}_{-\infty}d\tau_{1}\Big\langle P_{+}(t)H_{\text{int},R}(\tau_{3})H_{\text{int},R}(\tau_{2})H_{\text{int},L}(\tau_{4})H_{\text{int},L}(\tau_{1})\Big\rangle$
$=(-i)^{4}\text{Re}\int^{\infty}_{-\infty}dt_{4}\int^{\infty}_{0}dt_{3}\int^{\infty}_{0}dt_{2}\int^{\infty}_{0}dt_{1}\langle P_{+}\mathcal{G}(t_{4})d_{R}\mathcal{G}(t_{3})d_{R}\mathcal{G}(t_{2})d_{L}^{\dagger}\mathcal{G}(t_{1})d^{\dagger}_{L}\rangle$
$\qquad\langle E^{\dagger}(t-t_{4})E^{\dagger}(t-t_{4}-t_{3})E(t-t_{4}-t_{3}-t_{2})E(t-t_{4}-t_{3}-t_{2}-t_{1})\rangle$ (24)
In the expression using superoperator notation [37, 12], the variables $\tau_{j}$ denoted the exciton-photon interaction times in the first line. It is converted into time-delays $t_{j}=\tau_{ij}=\tau_{i}-\tau_{j}$ in the subsequent line. The field expression correlate the exciton evolution via the time integrals. Expanding the exciton correlation function in the basis of one and two-exciton states, and carrying out the implicit averaging over the phonons lead to an computable expression of the signal which contains the exciton Green’s functions (defined in Eq. II). We further recast the expression in the frequency domain to obtain
$O_{ff}(t)$ $=2\,\text{Re}\Big\{\int\frac{d\omega_{4}}{2\pi}\int\frac{d\omega_{3}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}}{2\pi}e^{+i\omega_{3}\tau_{3}+i\omega_{4}\tau_{4}-i\omega_{2}\tau_{2}-i\omega_{1}\tau_{1}}\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
$\Big\{d_{fe^{\prime}}d_{e^{\prime}g}d_{fe}^{*}d_{eg}^{*}\times\mathcal{G}_{ff}(-\omega_{4}-\omega_{3}+\omega_{2}+\omega_{1})\mathcal{G}_{fe^{\prime}}(-\omega_{3}+\omega_{2}+\omega_{1})\mathcal{G}_{fg}(\omega_{2}+\omega_{1})\mathcal{G}_{eg}(+\omega_{1})$
$+d_{fe^{\prime\prime}}d_{fe^{\prime\prime}}^{*}d_{eg}d_{eg}^{*}\times\mathcal{G}_{ff}(-\omega_{4}+\omega_{2}-\omega_{3}+\omega_{1})\mathcal{G}_{fe^{\prime\prime}}(+\omega_{2}-\omega_{3}+\omega_{1})\mathcal{G}_{e^{\prime\prime}e^{\prime\prime},ee}(-\omega_{3}+\omega_{1})\mathcal{G}_{eg}(+\omega_{1})$
$+d_{fe^{\prime}}d_{fe}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\mathcal{G}_{ff}(-\omega_{4}+\omega_{2}-\omega_{3}+\omega_{1})\mathcal{G}_{fe^{\prime}}(+\omega_{2}-\omega_{3}+\omega_{1})\mathcal{G}_{ee^{\prime}}(-\omega_{3}+\omega_{1})\mathcal{G}_{eg}(+\omega_{1})$
$+d_{fe^{\prime\prime}}^{*}d_{fe^{\prime\prime}}d_{eg}d_{eg}^{*}\times\mathcal{G}_{ff}(+\omega_{2}-\omega_{4}-\omega_{3}+\omega_{1})\mathcal{G}_{e^{\prime\prime}f}(-\omega_{4}-\omega_{3}+\omega_{1})\mathcal{G}_{e^{\prime\prime}e^{\prime\prime},ee}(-\omega_{3}+\omega_{1})\mathcal{G}_{eg}(+\omega_{1})$
$+d_{fe}^{*}d_{fe^{\prime}}d_{e^{\prime}g}d_{eg}^{*}\times\mathcal{G}_{ff}(+\omega_{2}-\omega_{4}-\omega_{3}+\omega_{1})\mathcal{G}_{ef}(-\omega_{4}-\omega_{3}+\omega_{1})\mathcal{G}_{e^{\prime}e}(-\omega_{3}+\omega_{1})\mathcal{G}_{eg}(+\omega_{1})\Big\}\Big\}$ (25)
In expression above is significantly more detailed. Here five types of phonon-averaged exciton Green’s functions, representing one, and two-exciton inter-manifold coherence, one and two-exciton transport and one-exciton intra-manifold coherence appear. They also bear accumulated frequencies of the driving field in their arguments. The first term represents the two-exciton coherence-mediated pathways, the second and fourth terms represents one-exciton transport-mediated pathways and the third and fifth terms represents the one-exciton coherence-mediated pathways. A more explicit form of the expression can be obtained if we plug in the expressions of the exciton Green’s functions. It gives
$O_{ff}(t)$ $\propto 2\mathrm{Re}\Big\{\int\frac{d\omega_{4}}{2\pi}\int\frac{d\omega_{3}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}}{2\pi}e^{+i\omega_{3}\tau_{3}+i\omega_{4}\tau_{4}-i\omega_{2}\tau_{2}-i\omega_{1}\tau_{1}}\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
$\Big\{\frac{d_{fe^{\prime}}}{-\omega_{4}-\omega_{3}+\omega_{2}+\omega_{1}-z_{ff}}\frac{d_{e^{\prime}g}}{-\omega_{3}+\omega_{2}+\omega_{1}-z_{fe^{\prime}}}\frac{d_{fe}^{*}}{+\omega_{2}+\omega_{1}-z_{fg}}\frac{d_{eg}^{*}}{\omega_{1}-z_{eg}}$
$+\frac{d_{fe^{\prime\prime}}}{-\omega_{4}+\omega_{2}-\omega_{3}+\omega_{1}-z_{ff}}\frac{d_{fe^{\prime\prime}}^{*}}{\omega_{2}-\omega_{3}+\omega_{1}-z_{fe^{{}^{\prime\prime}}}}$
$d_{e^{\prime}g}\left(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}\frac{1}{-\omega_{3}+\omega_{1}-z_{p}}\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}\frac{1}{-\omega_{3}+\omega_{1}-z_{ee^{\prime}}}\right)\frac{d_{eg}^{*}}{\omega_{1}-z_{eg}}$
$+\frac{d_{fe^{\prime\prime}}^{*}}{+\omega_{2}-\omega_{4}-\omega_{3}+\omega_{1}-z_{ff}}\frac{d_{fe^{\prime\prime}}}{-\omega_{4}-\omega_{3}+\omega_{1}-z_{fe^{\prime\prime}}}$
$d_{e^{\prime}g}\left(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}\frac{1}{-\omega_{3}+\omega_{1}-z_{p}}\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}\frac{1}{-\omega_{3}+\omega_{1}-z_{ee^{\prime}}}\right)\frac{d_{eg}^{*}}{\omega_{1}-z_{eg}}\Big\}\Big\}$ (26)
A compact version of this expression, using the abbreviation $I(\omega_{j};z_{ab})=(\omega_{j}+z_{ab})^{-1}$ is presented in Eqn. III.1. Proceeding via tedious yet straightforward frequency integrations yields the expression below
$O_{ff}(t)$ $\propto 2\text{Re}\Big[\exp{(z_{ff}t)}\Big(d_{fe^{\prime}}d_{e^{\prime}g}d_{fe}d_{eg}^{*}\times\big\langle E^{\dagger}(z_{fe^{\prime}}-z_{ff})E^{\dagger}(z_{fg}-z_{fe^{\prime}})E(-z_{eg}+z_{fg})E(z_{eg})\big\rangle$
$+d_{fe^{\prime\prime}}d_{fe^{\prime\prime}}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\big\langle E^{\dagger}(+z_{fe^{\prime\prime}}-z_{ff})E(-z_{p}+z_{fe^{\prime\prime}})\sum_{p}\chi^{R}_{e^{\prime\prime}p}D_{pp}^{-1}E^{\dagger}(z_{eg}-z_{p})\chi^{L}_{pe}E(z_{eg})\big\rangle$
$+d_{fe^{\prime}}d_{fe^{\prime}}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\big\langle E^{\dagger}(+z_{fe^{\prime}}-z_{ff})E(-z_{ee^{\prime}}+z_{fe^{\prime}})E^{\dagger}(z_{eg}-z_{ee^{\prime}})E(z_{eg})\big\rangle$
$+d_{fe^{\prime\prime}}d_{fe^{\prime\prime}}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\big\langle E^{\dagger}(-z_{e^{\prime\prime}f}+z_{ff})E(z_{p}-z_{e^{\prime\prime}f})\sum_{p}\chi^{R}_{e^{\prime\prime}p}D_{pp}^{-1}E^{\dagger}(z_{ee^{\prime}}-z_{p})\chi^{L}_{pe}E(z_{eg})\big\rangle$
$+d_{fe^{\prime\prime}}d_{fe^{\prime\prime}}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\big\langle E^{\dagger}(-z_{ef}+z_{ff})E(+z_{ee^{\prime}}-z_{ef})E^{\dagger}(z_{eg}-z_{ee^{\prime}})E(z_{eg})\big\rangle\Big)\Big]$ (27)
which is amenable to numerical simulation.
Appendix C Entangled photon correlation function
Here we present the numerical simulation of the square of the joint spectral amplitude function, for typical parameter regime used in the main simulation.
Figure 8: The squared joint spectral amplitudes of the entangled photon pairs are displayed, illustrating variations in the entanglement time parameter ($\tilde{T}_{\text{ent}}$) and the temporal width of the SPDC pump ($\tau_{0}$). They have noticeable implications for the nature of the frequency (anti-)correlations between the signal and idler pair, a feature that can be utilized for spectroscopic applications.
Appendix D Temporal spectral filtering functions
We use the following temporal ${F}_{\bar{t}}(t)$, and spectral $F_{\bar{\omega}}(\omega)$ filtering functions, expressed as
$F_{\bar{t}}(t^{\prime})$ $=\theta(t^{\prime}-\bar{t})\exp({-\sigma_{\bar{t}}(t^{\prime}-\bar{t})})$ (28)
$F_{\bar{\omega}}(\omega^{\prime})$ $=i(\omega^{\prime}-\bar{\omega}+i\sigma_{\omega})^{-1}$ (29)
These functions have four tunable parameters: the centering variables, $\bar{t}$, $\bar{\omega}$, and the width $\sigma_{\bar{t}}$, $\sigma_{\bar{\omega}}$. The presented simulations in Fig. 7 illustrates their role in signal classification.
Appendix E Note on the observable: photon coincidence signal
Here, we present a brief outline of derivation of the bare photon coincidence counting observable (i.e., without the filtering stage), following the original theory developed in [13, 11]. The expression is already presented in Eq. III, and appears as
$\langle N_{s}(t_{1}^{\prime},\tau_{1})N_{i}(t_{2}^{\prime},\tau_{2}^{\prime})\rangle=\int_{-\infty}^{t_{1}^{\prime}}dt_{1}\int_{-\infty}^{t_{1}^{\prime}+\tau_{1}}dt_{3}$
$\times\int_{-\infty}^{t_{2}^{\prime}}dt_{2}\int_{-\infty}^{t_{2}^{\prime}+\tau_{2}}dt_{4}\langle V_{R}^{\dagger}(t_{4})V_{R}^{\dagger}(t_{3})V_{L}(t_{1})V_{L}(t_{2})\rangle$
$\times\sum_{s,s^{\prime}}\sum_{i,i^{\prime}}\langle E_{s^{\prime}R}(t_{4})E_{i^{\prime}R}(t_{3})E_{s^{\prime}R}^{\dagger}(t_{2}^{\prime}+\tau_{2})E_{i^{\prime}R}^{\dagger}(t_{1}^{\prime}+\tau_{1})$
$\times E_{iL}(t_{1}^{\prime})E_{sL}(t_{2}^{\prime})E_{iL}^{\dagger}(t_{1})E_{sL}^{\dagger}(t_{2})\rangle.$ (30)
Here, the averaging of the field correlation term is carried out over the detector modes, and the auxiliary photon modes, where as the averaging of the exciton correlation term is carried out over two-exciton state following transport. In the presented theory, the leading order contributions to the signal arise from expansion of the time dependent density operator to the fourth order in emitted field. Assuming photon emission into vacuum modes, and following the standard algebraic manipulations one obtains
$\langle N_{s}(t_{1}^{\prime},\tau_{1})N_{i}(t_{2}^{\prime},\tau_{2}^{\prime})\rangle=\mathcal{D}^{2}(\omega_{1})\mathcal{D}^{2}(\omega_{2})$
$\times\langle V_{R}^{\dagger}(t_{2}^{\prime}+\tau_{2})V_{R}^{\dagger}(t_{1}^{\prime}+\tau_{1})V_{L}(t_{1}^{\prime})V_{L}(t_{2}^{\prime})\rangle.$ (31)
This expression represents the exciton correlation function weighted by the density of detector modes. The correlation function can be further expanded in terms of individual pathways Fig. 2, yielding the expressions in Eq. III.3.

### v2

@@
 We propose a protocol that combines photon-entanglement-enhanced narrowband excitation of two-exciton states with time-frequency-filtered two-photon coincidence counting.
-It can alleviate bottlenecks associated with probing exciton dynamics spread across multiple spectral and temporal windows.
-We demonstrate that non-classical correlations of entangled photon pairs can be used to prepare narrowband two-exciton population distributions, circumventing transport in mediating states.
-The distributions thus created can be monitored using time-frequency-filtered photon coincidence counting, and the pathways contributing to photon emission events can be classified by tuning filtering parameters.
+This approach alleviates bottlenecks associated with probing exciton dynamics spread across multiple spectral and temporal windows.
+We demonstrate that non-classical correlations of entangled photon pairs can be used to prepare narrowband two-exciton population distributions, thereby circumventing transport in mediating one-exciton states.
+The evolution of these population distributions and cascading transitions can be monitored using time-frequency-filtered photon coincidence counting.
 Numerical simulations for a light-harvesting aggregate highlight the ability of this protocol to achieve selectivity by suppressing or amplifying specific pathways.
-Combining entangled photonic sources and multidimensional photon counting allow promising applications to spectroscopy and sensing.
+Combining entangled photonic sources with multidimensional photon counting allows promising applications in spectroscopy and sensing.
 I Introduction
-In photosynthetic systems, light-harvesting complexes are critical components that facilitate efficient photon energy conversion through exciton formation and transport.
-Investigating the structural and functional mechanisms of exciton dynamics provides essential insights for engineering biomimetic energy-harvesting systems.
-In these systems, aggregated chromophores serve as the energy harvesters.
-The aggregation of individual chromophores results in delocalized electronic excitations: molecular excitons which are primarily of the Frenkel type.
-The excitons, often generated through photo-excitation, participate in multiscale relaxation and dephasing kinetics, due to the interaction with phonons [16, 1, 30], ultimately leading to the primary charge separation processes.
-Previous investigations have sought to monitor the role of coherent energy transfer that results from transition dipole-dipole interactions.
-The later can be probed via multi-pulse four-wave mixing techniques, popularly known as two-dimensional spectroscopy.
-The investigation multi-exciton states, where signatures of exciton-exciton correlations can be found, have comparatively received less attention.
-The early attributes of such correlations can already be seen in the two-exciton state energetics.
-Prior investigations have demonstrated that various modalities of two-photon absorption spectroscopy, resonant and off-resonant, can spectroscopically monitor these resonances [50, 51, 20, 15, 35, 54, 49, 7, 53, 26, 25, 33, 48, 21, 45, 29, 4].
-Probing state-resolved dynamics in the two-exciton manifolds is complicated by the high density of states that take part in phonon-induced dissipation spanning multiple time and energy scales.
-Over the past several years, a series of proposals, limited to atomic, and simple molecular systems have underscored the advantages of employing entangled photon sources to improve accessibility of doubly excited states while scaling favorably with the probe intensity [46, 23, 44, 40, 14, 22, 8, 18, 19, 17, 43, 52, 32, 5, 34].
-These proposals were complemented by examinations of the role of photon entanglement in the context of the reported enhancements [42, 28, 27, 39, 41, 24, 3, 6].
-Parallel theoretical investigations have established that exploiting the non-classical time-frequency correlations of photonic states may yield superior signal resolution in spectroscopic measurements compared to classical counterparts [10, 8].
-To monitor the fate of two-exciton states, multiphoton fluorescence intensity correlation measurements have been proposed [34, 31, 38, 11, 13].
-In a series of work [11, 13], the concept of multidimensional signal measurements was extended to include coincidence counting measurements.
-This theoretical framework promises to transparently link information derived from photon correlation measurements to molecular events.
-The time and frequency filtered $n$-photon coincidence counting measurements provide a $2n$-dimensional parameter space for the events to be classified.
-However to firmly establish their role as a routine spectroscopic tool in extended molecular systems, it requires a protocol combining them with efficient initial state preparation strategies.
-The two-exciton states in molecular aggregates, such as light-harvesting complexes, are typically prepared via the manifold of one-exciton states, which are subject to phonon-induced transport and dephasing.
-Furthermore, as the number of chromophoric sites increases, the number of one-exciton and two-exciton states increases as well.
-As a consequence, the number of interfering optical transitions may increase dramatically.
-Additionally, phonon-induced transport redistributes the population among the participating states during the dynamics.
-In this work, we demonstrate that state-selective entangled photon excitation and time-frequency-filtered multidimensional correlation spectroscopy can serve as an efficient probe for dissipative exciton kinetics.
-Specifically, this approach helps unravel two-exciton transport kinetics by monitoring cascaded photon emission events, tracking the correlation between these events, and benchmarking the nature of the exciton wavepacket created by entangled photon excitation.
+In interacting molecular aggregates, optical probing of dissipative exciton kinetics and the extraction of state-resolved spectroscopic information are inherently difficult.
+The high density of exciton states and exciton-phonon interactions lead to dissipative dynamics that span multiple time and energy scales.
+Molecular aggregates are found in naturally occurring photosynthetic systems.
+Within these photosynthetic systems, light-harvesting complexes are critical components that facilitate efficient photon energy conversion through exciton formation and transport.
+Investigating the structural and functional mechanisms of exciton dynamics provides essential insights for engineering biomimetic light-harvesting systems.
+[19, 1, 32].
+In these systems, the aggregation of individual chromophores results in delocalized electronic excitations.
+These molecular excitons are primarily of the Frenkel type.
+The generated excitons participate in relaxation and dephasing, leading to primary charge-separation processes.
+Signature of exciton-exciton interactions and correlations, which are typically hidden in the dynamics of multi-exciton states were investigated [38, 35, 53].
+The early attributes of such correlations can already be seen in the energy distributions of two-exciton states, which deviate from the sum of two one-exciton energies.
+Prior studies have demonstrated that various modalities of two-photon absorption spectroscopy can monitor the associated resonances.
+[54, 55, 23, 57, 29, 58, 8].
+In recent years, several proposals restricted to atomic and simple molecular systems have underscored the advantages of employing entangled photon pairs to probe two-exciton states [51, 25, 49, 44, 18, 24, 9, 21, 22, 20, 48, 56, 34, 4, 6].
+These proposals were complemented by critical examinations of the role of photon entanglement in enhancing two-photon absorption [47, 31, 30, 43, 46, 26, 2, 7].
+Theoretical investigations have also explored the role of non-classical time-frequency correlations in obtaining superior resolution in condensed-phase spectroscopies.
+[14, 9].
+Alongside, previous studies have proposed novel two-photon fluorescence intensity correlation measurements that can monitor the properties of two-exciton states [36, 33, 42, 15, 17, 50].
+In a series of theoretical works, time-frequency filtered photon coincidence counting was proposed as a technique that may provide spectroscopic information similar to that of two-dimensional spectroscopy, but at different timescales [15, 17].
+The optical excitation of two-exciton states in molecular aggregates, such as light-harvesting complexes, typically occurs via a manifold of one-exciton states subjected to population transport and dephasing.
+As a result, the population is redistributed among several states which may not be have highest propensity of transitions to the target two-exciton states.
+As the number of chromophoric sites increases, the number of one-exciton and two-exciton states increases as well.
+Consequently, the number of interfering optical transitions increase dramatically.
+The final two-exciton population also undergoes transport, dispersing any initially created narrowband two-exciton distribution.
+In this work, we demonstrate that two-photon excitation using entangled photon pairs may achieve the creation of narrow-band population distribution.
+We also show that the time-frequency-filtered two-photon counting can characterize cascaded photon emission events, resulting two-dimensional correlation plots.
+It constitutes a spectroscopic technique that can serve as an efficient probe for dissipative exciton dynamics.
 We focus on the light-harvesting complex LHCII, the most abundant antenna complex in plants and an important constituent of Photosystem II.
-It exists as a trimeric structure composed of structurally rigid monomeric aggregates; each monomer contains of several chlorophylls whose lowest energy levels are spectroscopically accessible.
-Exciton dynamics in LHCII exhibit significant complexity, particularly regarding the behavior of higher-order excitations.
-The combined protocol is shown to provide mechanistic insight into the cooperative exciton kinetics of LHCII, even in the presence of spectral weight redistribution caused by mode mixing and relaxation.
-Figure 1: (A) Illustration of the proposed protocol, consisting of the driving, evolution, and detection stages.
-The process is assisted by entangled photon pairs and temporal-spectral filtering of emitted photons.
-In (B), three plausible snapshots of the dynamics (I)-(III), featuring increasingly complex pathways, that may be monitored using the protocol is displayed.
-In the simulation, the exciton transport within both manifolds is accounted for.
-Since these pathways may be simultaneously present in the dynamics, the protocol allows for selective probing and aids in constructing a mechanistic narrative.
-In the following, in section II, we introduce the exciton-phonon Hamiltonian of the light-harvesting complex LHCII, which serves as our test system, and the photonic sources.
+Figure 1: (A) Illustration of the proposed protocol, classified as a three-stage process: preparation of the two-exciton population, dissipative evolution of the two-exciton population, and time-frequency filtered detection of cascaded photon emission.
+The action of the time filter (governed by $F_{\bar{t}}^{(j)}$) precedes the action of the frequency filter (governed by $F_{\bar{\omega}}^{(j)}$) in both cases.
+The vertical straight arrows denote interaction with the driving photon modes, and the wiggly arrows denote interaction with the photon modes of the detector.
+In (B), three exemplery dynamical pathways are shown.
+The tilted double-ended arrows denote intra-manifold exciton transport.
+In (I), both two- and one-exciton transport preserves the narrowband nature of the created population distribution.
+In (II), during the excitation stage, the participation of one-exciton transport alters the target two-exciton population distribution.
+The population distribution disperses due to two-exciton transport, which results in broadband cascaded emissions.
+In (III), the narrowband two-exciton population distribution funnels to an even narrower distribution, but the subsequent one-exciton transport erases the memory.
+It lead leads to a narrowband emission followed by a broadband emission.
+The filtering functions assist in constructing a mechanistic narrative by classifying these pathways.
+It exists as a trimeric structure composed of structurally rigid monomeric aggregates; each monomer contains several chlorophylls whose lowest three energy levels are spectroscopically accessible.
+In Section II, we introduce the exciton-phonon Hamiltonian of the light-harvesting complex LHCII.
+We also the exciton-photon Hamiltonian.
 This is complemented by the Appendix A.
-In the subsequent three subsections of Sec.
-III (III.1III.3), we detail the theoretical framework and numerical simulations of the combined protocol.
+In the subsequent sections, namely Sec.
+III (III.1III.3), we detail the theoretical framework for the combined protocol and numerical simulations.
 The manuscript concludes in Sec.
-IV with a summary of our findings, a discussion of future outlook, and a brief examination of the current limitations.
+IV with a summary of our findings and a brief commentary of the current limitations.
 II Driven dissipative one- and two-exciton kinetics
-The dissipative one- and two-exciton dynamics of the light-harvesting complex LHCII are simulated using the Frenkel exciton model within the Heitler-London approximation.
-The model is constructed by considering the lowest three electronic energy states of each of the chlorophyll molecules.
+We simulate the one- and two-exciton dynamics of the light-harvesting complex LHCII using the Frenkel exciton model within the Heitler-London approximation.
+The model is constructed by considering the lowest three electronic energy states for each chromophoric site.
 The Hamiltonian (with $\hbar=1$) is given by,
@@
 $+\sum_{j}\omega_{j}(b^{\dagger}_{j}b_{j}+\frac{1}{2})+\sum_{m,j}g_{m,j}(b^{\dagger}_{j}+b_{j})B_{m}^{{\dagger}}B_{m}$ (1)
-Here, the on-site excitation energy, $E_{m}$, and inter-site hopping, $J_{mn}$, are accounted for by the first term in the Hamiltonian.
-The Coulomb interaction, responsible for the inter-site hopping, results in the delocalization of the exciton states across the sites.
-The higher-order excitonic nonlinearities that shift exciton energy states are given by $U_{m}^{(1)}$ and $U_{mn}^{(2)}$.
-The exciton creation (annihilation) operators, $B_{m}^{\dagger}$ ($B_{m}$), follow the commutation relation $B_{n},B_{m}^{\dagger}]=\delta_{mn}(1-\eta(B_{m}^{\dagger}B_{m})^{2})$, where $\eta=3/2$, while the phonon creation (annihilation) operators, $b_{j}^{\dagger}$ ($b_{j}$), follow $b_{i},b_{j}^{\dagger}]=\delta_{ij}$.
-The two-exciton Hamiltonian can be constructed as $H_{mnkl}^{(2)}=H_{mn}^{(1)}\delta_{kl}+\delta_{mn}H_{kl}^{(1)}$, where $H_{mn}^{(1)}$ is given by the first line of the Eq.
-II.
+where the exciton creation (annihilation) operators, $B_{m}^{\dagger}$ ($B_{m}$), follow the commutation relation $B_{n},B_{m}^{\dagger}]=\delta_{mn}(1-\eta(B_{m}^{\dagger}B_{m})^{2})$ (with $\eta=3/2$) and the phonon creation (annihilation) operators, $b_{j}^{\dagger}$ ($b_{j}$), follow the commutation relation $b_{i},b_{j}^{\dagger}]=\delta_{ij}$.
+In the above, $E_{m}$ and $J_{mn}$ describes the on-site excitation and inter-site hopping, respectively.
+The later arises due the Coulomb interaction and leads to the delocalization of the exciton states across the sites.
+The one-exciton Hamiltonian can be constructed by considering the first term, i.e., $h_{mn}^{(1)}=(E_{m}\delta_{mn}+J_{mn})B_{m}^{\dagger}B_{n}$.
+Upon diagonalization, we obtain the corresponding one-exciton eigenstates $|e_{j}\rangle$.
+The two-exciton Hamiltonian can be constructed by considering $h_{mnkl}^{(2)}=h_{mn}^{(1)}\delta_{kl}+\delta_{mn}h_{kl}^{(1)}+U_{m}^{(1)}\delta_{mk}\delta_{nl}\delta_{mn}+U_{mn}^{(2)}\delta_{mk}\delta_{nl}$ where $h_{mn}^{(1)}$ is given above.
+We note that the matrix elements of the two-exciton Hamiltonian differs from the sum of two one-exciton Hamiltonian by the last two terms.
+These terms which contribute to two-exciton nonlinearities, represent the energetic shift of overtone and combination exciton states respectively.
+In a similar manner as the previous case, the two-exciton eigenstates $|f_{k}\rangle$ are obtained by diagonalization.
+The parameters for the Frenkel exciton Hamiltonian were taken from [41, 40, 10].
 For the LHCII, the number of chromophoric sites is $N_{s}=14$, which yields $N_{g}=1$ ground state, $N_{e}=14$ one-exciton states, and $N_{f}=105$ two-exciton states, composed of $14$ overtone and $91$ combination states.
-The excitonic eigenstates are obtained by the exact diagonalization of the exciton-number conserving Hamiltonian blocks as:
+The eigenstates can be expressed as
 $|e_{j}\rangle$ $=\sum_{m}T^{(1)}_{j,m}B_{m}^{\dagger}|0\rangle$
 $|f_{k}\rangle$ $=\sum_{mn}T^{(2)}_{k,mn}B_{m}^{\dagger}B_{n}^{\dagger}|0\rangle$ (2)
-These transformation coefficients, $T^{(n)}$ (for the nthn-th manifold) define the spectral weights associated with the delocalized eigenstates.
-It is noteworthy that the energies of the two-exciton overtone and combination states, dependent on the nonlinearity terms $U_{m}^{(1)}$ and $U_{mn}^{(2)}$ respectively, deviate from the sum of two one-exciton states.
-As a consequence, the relative ordering of states within the two-exciton manifold is contingent upon two competing sets of parameters, governing exciton delocalization and localization.
-The latter influence, via excited state absorption processes, all three pathways responsible for the two-photon absorption events addressed in this article.
-The Green’s functions (valid for both one and two exciton manifold) describing can be represented in a compact form as follows
+where $T^{(1)}_{j,m}$ and $T^{(2)}_{k,mn}$ denote the matrix elements of the diagonalization matrices mentioned above.
+These matrix elements determine the spectral weights associated with the eigenstates which are delocalized over the site basis.
+Figure 2: One- and two-exciton energies are displayed in the left and right panels, respectively.
+Nonuniform energy distributions are visible in both manifolds.
+Simulations in Section III.1 refer to the excitation of two-exciton states $f_{07}$ ($E_{f_{07}}=30192.9\,\text{cm}^{-1}$) and $f_{83}$ ($E_{f_{83}}=31243.2\,\text{cm}^{-1}$) via $e_{07}$ ($E_{e_{07}}=15307.1\,\text{cm}^{-1}$) and $e_{09}$ ($E_{e_{09}}=15655.9\,\text{cm}^{-1}$).
+The one exciton manifold spans from $E_{e_{1}}=14964.5\,\text{cm}^{-1}$ to $\omega_{e_{14}}=15934.7\,\text{cm}^{-1}$, hosting $14$ states within a spectral width of $\approx 970\,\text{cm}^{-1}$.
+The two-exciton manifold spans from $E_{f_{1}}=29844.3\,\text{cm}^{-1}$ to $E_{f_{105}}=31811.8\,\text{cm}^{-1}$, hosting $105$ states within a spectral width of $\approx 1968\,\text{cm}^{-1}$.
+The corresponding energy distributions, plotted in Fig.
+2, offers a visualization of the nature of the density of states in each manifold.
+We note that the energetic ordering of exciton eigenstates is dependent upon two competing parameters: $J_{mn}$ that govern exciton delocalization and $U_{m}^{(1)}$, $U_{mn}^{(2)}$ that govern exciton nonlinearities.
+The dissipative exciton dynamics is described by the Liouville-space Green’s function formalism. Using the exciton eigenbasis, we can express the Green’s function as follows
 $\mathcal{G}_{a_{1}a_{2},a_{3}a_{4}}(t)=\delta_{a_{1}a_{2}}\delta_{a_{3}a_{4}}\theta(t)[\exp{(-Kt)}]_{a_{1}a_{1},a_{3}a_{3}}$
 $+(1-\delta_{a_{1}a_{2}})\delta_{a_{1}a_{3}}\delta_{a_{2}a_{4}}\theta(t)\exp{(-i\omega_{a_{1}a_{3}}t-\gamma_{a_{1}a_{3}}t)}$ (3)
-The optical properties of exciton aggregates consisting of multiple three level emitters are primarily dictated by the nature of intermolecular interactions and two-exciton nonlinearities [26, 2, 1].
-The interaction with photonic sources, within the rotating wave approximation, can be described by the following Hamiltonian
-$H_{\mathrm{int}}(t)=EV^{\dagger}+E^{\dagger}V=\sum_{j,pq\in\{ge,ef\}}\sqrt{\frac{2\pi\omega_{j}}{V}}$
+The expression above is valid for both exciton transport and dephasing, which are captured by the first and the second terms, respectively. It is discussed further in Appendix A.
+Interaction with the photonic sources, within the rotating-wave approximation, can be described by the following Hamiltonian:
+$H_{\mathrm{int}}(t)=EV^{\dagger}+E^{\dagger}V=\sum_{j,pq\in\{ge,ef\}}\sqrt{\frac{2\pi\omega_{j}}{\Omega}}$
 $\Big(a_{j}\exp{(-i\omega_{j}t)}\exp{(i\omega_{pq}t)}B_{pq}+\mathrm{h.c.}\Big)$ (4)
-where $B_{pq}(t)=d_{pq}|p\rangle\langle q|(\text{with properties}\,\,\{p,q\}\in g,e,f$, and $a_{j}$ are the exciton and photon annihilation operators, describing inter-manifold exciton transition and excitations in the photonic modes, respectively.
-Figure 2: (A) Diagrammatic illustration (using the Schwinger-Keldysh loop representation) of the pathways involved in the proposed protocol, consisting of the driving, evolution, and emission stages.
-(B)-(C) Diagrammatic illustration (using the Liouville superoperator representation, suitable for dissipative systems) of the pathways involved in filtered emission (B) and excitation (C).
-The combinatorial pairing of five possible excitation diagrams and four possible emission diagrams gives rise to all possible pathways that contribute to the final signal.
-The corresponding key equations are presented in Section III.
-For our case, the deployed photonic sources are entangled photon pairs generated via spontaneous parametric down-conversion (SPDC) in the weak-down-conversion limit.
-In a typical scenario, a birefringent crystal is pumped with a classical pump that precedes to the generation of correlated photon pairs with frequencies $\omega_{s},\omega_{i}$.
-It involves two field-matter interactions with the incoming classical field modes and four interactions with the outgoing signal-idler fields.
-The correlation properties of the emitted photon field is determined by the physical properties of the SPDC crystal [44, 47, 23, 46, 12, 8, 9].
-The incoming entangled photon states can be described by the two photon state of the photon field
+In the above, we used a mode expansion of the photonic field operators, where $a_{j}$ denotes the photon annihilation operators and $\omega_{j}$ represents the mode frequencies.
+These mode operators follow bosonic commutation relations.
+The exciton annihilation operators, expressed in the eigenbasis, are denoted as $B_{pq}(t)=d_{pq}|p\rangle\langle q|(\text{with properties}\,\,\{p,q\}\in g,e,f$.
+The entangled photon pairs induce inter-manifold exciton transitions.
+Figure 3: (A) The Schwinger-Keldysh loop diagram describing the pathways involved in the proposed protocol.
+The excitation and emission states involve four exciton-photon interactions each with the driving and detection photon modes.
+The Liouville space ladder diagrams in (B)-(C) are suitable for capturing the processes in dissipative systems.
+The excitation and emission pathways are depicted in (B) and (C) respectively.
+In the excitation stage, described in (C), diagrams (II) and (IV) depends on one-exciton transport, $|e\rangle\langle e|\rightarrow|e^{\prime}\rangle\langle e^{\prime}|$.
+Diagram (I) depends on inter-manifold two-exciton coherence, $|f\rangle\langle g|$, while (III) and (V) depend on intra-manifold one-exciton coherence, $|e\rangle\langle e^{\prime\prime}|$.
+In the emission stage, described in (B), all diagrams have the two-exciton population distribution following transport as the initial state, $|f\rangle\langle f|\rightarrow|f^{\prime}\rangle\langle f^{\prime}|$.
+A combinatorial pairing of five excitation and four possible emission diagrams generate all possible pathways that may contribute to the signal.
+Their relative contributions to the final signal is jointly determined by the four-point photon correlation functions and the filtering properties.
+The corresponding equations are presented in Section III.
+We consider entangled photon pairs generated via spontaneous parametric down-conversion (SPDC) in the weak-down-conversion limit.
+In a typical scenario, a birefringent crystal is pumped with a classical laser, leading to the generation of correlated photon pairs with frequencies $\omega_{1}$ and $\omega_{2}$.
+The SPDC process involves two photon-matter interactions with the incoming classical field modes and four photon-matter interactions with the outgoing signal-idler fields.
+The correlation properties of the emitted photon field are determined by the physical properties of the SPDC crystal.
+[49, 52, 25, 51, 16, 9, 10].
+The generated entangled two-photon states can be described as follows:
 $\rho_{\text{photon}}$ $=\int\frac{d\omega_{1}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}^{\prime}}{2\pi}\int\frac{d\omega_{2}^{\prime}}{2\pi}$
 $F(\omega_{1},\omega_{2})F^{*}(\omega_{1}^{\prime},\omega_{2}^{\prime})a_{\omega_{1}}^{{\dagger}}a_{\omega_{2}}^{{\dagger}}|0\rangle\langle 0|a_{\omega_{1}^{\prime}}a_{\omega_{2}^{\prime}}$ (5)
-The entangled photon field correlation function is obtained as,
-$\langle E^{{\dagger}}(\omega_{b}^{\prime})E^{{\dagger}}(\omega_{a}^{\prime})E(\omega_{b})E(\omega_{b})\rangle=F^{*}(\omega_{b}^{\prime},\omega_{a}^{\prime})F(\omega_{a},\omega_{b})$ (6)
-where,
-$F(\omega_{a},\omega_{b})$ $=\alpha A_{p}(\omega_{p})\Big\{\mathrm{sinc}[\phi(\omega_{a},\omega_{b})]$
-$\exp{(i\phi(\omega_{a},\omega_{b}))}+\omega_{1}\leftrightarrow\omega_{2}\Big\}\text{with}$
-$\phi(\omega_{a},\omega_{b})$ $=(\omega_{a}-\omega_{1})\tilde{T}_{1}/2+(\omega_{b}-\omega_{1})\tilde{T}_{2}/2$ (7)
-where $\alpha$ governs the down-conversion efficiency.
-The temporal entanglement parameter of the photon pair sources is given by $\tilde{T}_{\mathrm{ent}}=\tilde{T}_{2}-\tilde{T}_{1}$ establish the upper bound for the interaction time interval.
-The transform-limited field profile of the SPDC pump, $A_{p}(\omega_{a},\omega_{b})$ is taken as a Gaussian, given by, $A_{p}(\omega_{a},\omega_{b})=E_{0}\sqrt{\pi/\Gamma_{p,0}}\exp{[-(\omega-\omega_{a}-\omega_{b})^{2}/4\Gamma_{p,0}]}$, where $\Gamma_{p,0}$ is the temporal width.
-The quantity is dependent on the time delays acquired by the signal and idler photons during the propagation within the nonlinear crystal.
-In contrast to entangled photon pairs, the correlation function of coherent laser fields would have factorized into the product of four amplitude functions, i.e.
-$\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
-$=A^{*}_{4}(\omega_{4})A^{*}_{3}(\omega_{3})A_{1}(\omega_{2})A_{1}(\omega_{1})$ (8)
-and the signal would have scaled with the square of the incoming field intensity.
+The four-point correlation function of the entangled photon pairs is relevant for the creation of two-exciton states. Using the expression above, it can be evaluated, yielding
+$\langle E^{{\dagger}}(\omega_{b}^{\prime})E^{{\dagger}}(\omega_{a}^{\prime})E(\omega_{b})E(\omega_{b})\rangle_{\rho_{\text{photon}}}$
+$=F^{*}(\omega_{b}^{\prime},\omega_{a}^{\prime})F(\omega_{a},\omega_{b})$ (6)
+We defined, $F(\omega_{a},\omega_{b})=\alpha A_{p}(\omega_{p})\Big\{\mathrm{sinc}[\phi(\omega_{a},\omega_{b})]$, where $\phi(\omega_{a},\omega_{b})=(\omega_{a}-\omega_{1})\tilde{T}_{1}/2+(\omega_{b}-\omega_{1})\tilde{T}_{2}/2$ and $\alpha$ denotes the parameter governing the down-conversion efficiency.
+The entanglement time parameter $\tilde{T}_{\mathrm{ent}}=\tilde{T}_{2}-\tilde{T}_{1}$ establishes an upper bound for the intervals between two photon-exciton interactions and depends on the time delays acquired by the signal and idler photons during propagation within the nonlinear crystal.
+It characterizes the non-classical correlation in the photonic source.
+The classical transform-limited pulse driving the SPDC crystal is defined as follows: $A_{p}(\omega_{a},\omega_{b})=E_{0}\sqrt{\pi/\Gamma_{p,0}}\exp{[-(\omega-\omega_{a}-\omega_{b})^{2}/4\Gamma_{p,0}]}$, where $\Gamma_{p,0}$ is the temporal width.
 III Results
-The excitation-evolution-emission events occur in three stages.
-Excitation by entangled photon pairs creates a tailored two-exciton wavepacket.
-The two-exciton wavepacket or density operator populations undergo transport, leading to a redistribution of populations over a finite waiting time.
-This redistributed population serves as the initial configuration for photon emission events.
-The cascaded two-photon emission process proceeds via two one-photon emissions interleaved by exciton transport in the one-exciton manifold.
-The signal can be expressed in a compact superoperator form
+The sequence of processes: entangled photon excitation, dissipative evolution, and two-photon emission occurs in three stages, as depicted in Fig.
+1 (A) (I) to (III).
+Excitation by entangled photon pairs creates a tailored two-exciton population distribution which undergoes transport for a finite time.
+It results in a redistribution of populations.
+The redistributed population serves as the initial configuration for the cascaded two-photon emission processes.
+The emission process involves two successive one-photon emission events interleaved by one-exciton transport in the one-exciton manifold.
+In Fig.
+1 (B), we demonstrate three plausible pathways via which the dynamics may proceed.
+Due to transport, the two-exciton population distribution may lose its narrowband character or may even become narrower.
+The dynamical information is encoded into the photon emissions that correspond to the inter-manifold transitions between the upper two manifolds ($|f\rangle\rightarrow|e\rangle$).
+The one-exciton transport that follows may also create additional dispersion of the one-exciton population.
+The corresponding dynamical information is encoded into the inter-manifold transition between lower two manifolds ($|e\rangle\rightarrow|g\rangle$).
+These two sets of photon emissions can be classified by controlling the properties of the time and frequency filters, denoted $F_{\bar{t}}^{(j)}$ and $F_{\bar{\omega}}^{(j)}$ in the Fig.
+1.
+In the subsequent sections, we use the following notation for Liouville-space operators: $A_{\alpha}$, where the subscript denotes $\alpha=L,R,+,$-.
+These operators are defined by their action on a Hilbert-space operator $X$ as $A_{L}X\equiv AX$ and $A_{R}X\equiv XA$.
+The matrix elements are obtained as $(A_{\alpha}X)_{ij}\equiv\sum_{kl}(A_{\alpha})_{ij,kl}X_{kl}$.
+We also define $A_{+}\equiv\frac{1}{2}(A_{L}+A_{R})$ and $A_{-}\equiv A_{L}-A_{R}$, where the matrix elements are given by: $(A_{-})_{ij,kl}=A_{ik}\delta_{jl}-A_{\ell j}\delta_{ik}$ and $(A_{+})_{ij,kl}=\frac{1}{2}[A_{ik}\delta_{jl}+A_{lj}\delta_{ik}]$.
+The time-frequency filtered two-photon coincidence signal can be expressed as follows:
 $S(\overline{t}_{s},\overline{\omega}_{s};\overline{t}_{i},\overline{\omega}_{i})$ $=\int_{-\infty}^{\infty}dt_{s}^{\prime}d\tau_{s}\int_{-\infty}^{\infty}dt_{i}^{\prime}d\tau_{i}\sum_{s,s^{\prime},i,i^{\prime}}\text{Tr}_{s,s^{\prime},i,i^{\prime}}[D^{(s,s^{\prime})}(\overline{t}_{s},\overline{\omega}_{s};t_{s}^{\prime},\tau_{s})D^{(i,i^{\prime})}(\overline{t}_{i},\overline{\omega}_{i};t_{i}^{\prime},\tau_{i})B^{(s,s^{\prime},i,i^{\prime})}(t_{s}^{\prime},\tau_{s};t_{i}^{\prime},\tau_{i})]$ (7)
-where the superoperator for filtered detection
+Here we define the superoperator for filtered detection:
 $D^{(\nu,\nu^{\prime})}(\overline{t}_{\nu},\overline{\omega}_{\nu};t_{\nu}^{\prime},\tau_{\nu})$
 $=D(\overline{t}_{\nu},\overline{\omega}_{\nu};t_{\nu}^{\prime},\tau_{\nu})E_{\nu^{\prime}L}(t_{\nu}^{\prime})E_{\nu R}^{\dagger}(t_{\nu}^{\prime}+\tau_{\nu})$ (8)
-and the bare two-photon signal
+and for the bare two-photon signal
 $B^{(s,s^{\prime},i,i^{\prime})}(t_{s}^{\prime},\tau_{s};t_{i}^{\prime},\tau_{i})$
 $=T_{L}^{(s^{\prime},i^{\prime})}(t_{s}^{\prime},t_{i}^{\prime})T_{R}^{(s,i)\dagger}(t_{s}^{\prime}+\tau_{s},t_{i}^{\prime}+\tau_{i})$ (9)
-is defined using the following superoperators corresponding to emission events
+We also use the following superoperators to define the emission processes:
 $T_{R}^{{(s^{\prime},i^{\prime})},{\dagger}}(t_{s}^{\prime},t_{i}^{\prime})$ $=i^{2}\int_{-\infty}^{t_{2}^{\prime}+\tau_{2}}dt_{4}\int_{-\infty}^{t_{1}^{\prime}+\tau_{1}}dt_{3}$
@@
 $E_{sL}^{\dagger}(t_{1})E_{iL}^{\dagger}(t_{2})V_{L}(t_{1})V_{L}(t_{2})$ (10)
-The averaging has been carried out over the two-exciton density operator following transport, i.e., $\rho_{f^{\prime}f^{\prime}}(t)=\mathcal{G}_{f^{\prime}f^{\prime},ff}(t-t^{\prime})\rho_{ff}(t^{\prime})$.
-Under a set of physically motivated approximations, the averaging can be carried out over each of these three terms independently.
-We also define the entangled photon driven excitation superoperator for future convenience, given by
-$O(t)=T_{R}^{{\dagger}}(\tau_{2},\tau_{3})T_{L}(\tau_{1},\tau_{4})=i^{2}\int_{-\infty}^{\tau_{2}^{\prime}}d\tau_{4}\int_{-\infty}^{\tau_{1}^{\prime}}d\tau_{3}$
+Here the superoperators $T_{L}$ and $T_{R}$ are defined by their action from the left and right sides of the density operator, which describes the initial configuration for the photon emission following the excitation and two-exciton transport, i.e., $\rho_{f^{\prime}f^{\prime}}(t)=\mathcal{G}_{f^{\prime}f^{\prime},ff}(t-t^{\prime})\rho_{ff}(t^{\prime})$.
+For future convenience, we also define the observable that corresponds to the preparation of the two-exciton population distribution
+$O(t)=\tilde{T}_{R}^{{\dagger}}(\tau_{2},\tau_{3})\tilde{T}_{L}(\tau_{1},\tau_{4})=i^{2}\int_{-\infty}^{\tau_{2}^{\prime}}d\tau_{4}\int_{-\infty}^{\tau_{1}^{\prime}}d\tau_{3}$
 $\int_{-\infty}^{\tau_{3}}d\tau_{2}\int_{-\infty}^{\tau_{4}}d\tau_{1}E_{L}^{\dagger}(\tau_{4})E_{L}^{\dagger}(\tau_{1})E_{R}(t_{3})E_{R}(\tau_{2})$
 $\qquad\qquad V_{R}^{{\dagger}}(\tau_{3})V_{R}^{{\dagger}}(\tau_{2})V_{L}(\tau_{4})V_{L}(\tau_{1})$ (11)
-In the above, we assume that the excitation and emission events are not temporally overlapping, which results in the signal in Eq.
-9 being factorized in terms of three principal contributions: first, preparation, involving exciton-photon (driving fields) interactions that lead to the preparation of the two-exciton density operator; second, evolution, describing two-exciton transport over a parametric time interval; and third, detection, describing exciton-photon (detection fields) interactions that yield the time-frequency filtered two-photon coincidence.
-In subsequent sections, we systematically describe all three contributions.
+Going forward, we assume that the excitation and emission stages are temporally non-overlapping, which results in the factorization of Eq.
+7 into separate contributions from each of the three stages.
+In the following sections, we systematically analyze the role of each stage.
 Figure 4: The two-exciton population distribution map, constructed by sequentially targeting each of the $105$ states in that manifold, using degenerate entangled photon pairs i.e., $\omega_{i}=\omega_{p}/2=E_{\mathrm{target}}^{(2)}/2$.
@@
 III.1 Entangled photon-pair excitation
-The quantity of interest at the preparation stage, the two-exciton population distribution, can be defined using a projection operator acting on the diagonal matrix elements of the phonon-averaged exciton density operator, expressed as: $\rho_{ff}(t)=\text{Tr}[O(t)|f\rangle\langle f|]_{\rho_{0}}$.
-Here, a factorizable initial condition between the driving, the detection field modes, and the exciton density operator is assumed.
-We perturbatively expand the combined photon-exciton-phonon density operator to the fourth order in $H_{\text{int}}(t)$, and partially trace over the phonon modes to obtain the desired expression for the observable (derivation detailed in Appendix B) as follows:
+The quantity of interest at the preparation stage, the two-exciton population distribution, is defined by the projection operator $|f\rangle\langle f|$ acting on the two-exciton density operator, expressed as $\rho_{ff}(t)=\text{Tr}[O(t)|f\rangle\langle f|]$, where the partial trace $\text{Tr}_{B}$ is taken over the phonon modes.
+We perturbatively expand the density operator to the fourth order in $H_{\text{int}}(t)$ to obtain the expression for the observable:
 $\rho_{ff}(t)$ $\propto 2\text{Re}\Big\{\int\frac{d\omega_{4}}{2\pi}\int\frac{d\omega_{3}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}}{2\pi}e^{+i\omega_{3}\tau_{3}+i\omega_{4}\tau_{4}-i\omega_{2}\tau_{2}-i\omega_{1}\tau_{1}}\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
@@
 $+d_{fe^{\prime}}d_{fe}^{*}d_{e^{\prime}g}d_{eg}^{*}\times I(-\omega_{4}+\omega_{2}-\omega_{3}+\omega_{1};-z_{ff})I(\omega_{2}-\omega_{3}+\omega_{1};-z_{fe^{{}^{\prime\prime}}})$
-$\left(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}I(-\omega_{3}+\omega_{1};-z_{p})\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}I(-\omega_{3}+\omega_{1};-z_{ee^{\prime}})\right)I(\omega_{1};-z_{eg})$
+$\Big(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}I(-\omega_{3}+\omega_{1};-z_{p})\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}I(-\omega_{3}+\omega_{1};-z_{ee^{\prime}})\Big)I(\omega_{1};-z_{eg})$
 $+d_{fe^{\prime}}d_{fe}^{*}d_{e^{\prime}g}d_{eg}^{*}\times I(+\omega_{2}-\omega_{4}-\omega_{3}+\omega_{1};-z_{ff})I(-\omega_{4}-\omega_{3}+\omega_{1};-z_{fe^{\prime\prime}})$
-$\left(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}I(-\omega_{3}+\omega_{1};-z_{p})\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}I(-\omega_{3}+\omega_{1};-z_{ee^{\prime}})\right)I(\omega_{1};-z_{eg})\Big\}\Big\}$ (14)
-It is a sum of five distinct terms that represent interfering exciton pathways (see Liouville space Feynman diagrams presented in Fig.
-2 for more insights).
-It includes contributions from both one- and two-exciton coherence and population terms of the exciton density operator.
-The first term represents the two-exciton coherence-mediated pathway, which is susceptible to inter-manifold dephasing.
-The second and fourth terms describe pathways susceptible to one-exciton population transport.
-The third and fifth terms describe pathways susceptible to intra-manifold dephasing.
-The convolutional form is particularly advantageous for introducing the field correlation functions, given in Eq.
-6-Eq.
+$\Big(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}I(-\omega_{3}+\omega_{1};-z_{p})\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}I(-\omega_{3}+\omega_{1};-z_{ee^{\prime}})\Big)I(\omega_{1};-z_{eg})\Big\}\Big\}$ (12)
+The derivation is outlined in Appendix C.
+The expression above is a sum of five distinct terms, each representing an exciton pathway that contributes to the signal.
+The Liouville-space Feynman diagrams in Fig.
+3 (I)–(V), offer further insights into the exciton processes captured by this expression.
+The pathways include contributions from both exciton coherences and populations.
+The first term represents the two-exciton coherence-mediated pathway (diagram (I)), which is susceptible to inter-manifold dephasing.
+The second and fourth terms describe pathways susceptible to one-exciton population transport (diagrams (II) and (IV)).
+The third and fifth terms describe pathways susceptible to intra-manifold dephasing (diagrams (III) and (V)).
+The convolutional form provided above is particularly advantageous for exploring the role of entangled photon pairs.
+Their effect is included via the four-point correlation functions, defined in Eq.
 II.
-The field correlation functions can be manipulated to selectively suppress and amplify specific exciton resonances.
-Next, we illustrate these capabilities via numerical simulations for both degenerate and nondegenerate entangled photon pairs.
+The following two sections use both degenerate and nondegenerate entangled photon pairs to illustrate these capabilities.
 III.1.1 Simulation: excitation of the full suite of two-exciton states using degenerate entangled photon pairs
-In this section, we monitor the final population distribution following the excitation via spectrally degenerate entangled photon pairs, i.e., $\omega_{1}=\omega_{2}=\omega_{p}/2=E_{\mathrm{target}}^{(2)}/2$.
-We had carried out a parametric scan, taking each of the $N_{f}=105$ two-exciton states as a target, and present the final normalized population distribution in Fig.
-3.
-The simulation presented in the left column uses an entanglement time of $T_{\mathrm{ent}}=10\,\mathrm{fs}$ and an SPDC pump width of $\tau_{0}=150\,\mathrm{fs}$, and acts as a reference case.
-Variations in the temporal width of the SPDC pump ($\tau_{0}=50\,\text{fs}$) and the entanglement time parameter ($T_{\text{ent}}=60\,\text{fs}$) are presented in the middle and the right column, respectively.
-We note that target states near the middle of the energy ladder exhibited broader distributions around the target in all three cases.
-In other words, we find a lower degree of selectivity for target states in the middle rung of the two-exciton manifold and higher selectivity for those belonging to the lower rung.
-Altering the width of SPDC pump pulse in the middle column resulted in diminished selectivity, likely due to the activation of multiple pathways.
-Altering the entanglement time in the right column preserved the selectivity for the majority of states.
+In this section, we monitor the final two-exciton population distribution following excitation by spectrally degenerate entangled photon pairs, i.e., $\omega_{1}=\omega_{2}=\omega_{p}/2=E_{\mathrm{target}}^{(2)}/2$.
+For an SPDC source pumped by a finite-bandwidth laser, the generated photon pairs usually carry a range of frequencies.
+The photon pairs, emitted within a specific solid angle, must be filtered to select the degenerate components.
+We carry out a parametric scan, taking each of the $N_{f}=105$ two-exciton states as a target, and present the final normalized population distribution in Fig.
+4.
+The results presented in the left column use an entanglement time of $T_{\mathrm{ent}}=10\,\mathrm{fs}$ and an SPDC pump width of $\tau_{0}=150\,\mathrm{fs}$, which acts as a reference case.
+The middle and right columns present simulations with a variation in the temporal width of the SPDC pump ($\tau_{0}=50\,\mathrm{fs}$) and the entanglement time parameter ($T_{\mathrm{ent}}=60\,\mathrm{fs}$), respectively, while keeping all other parameters unchanged.
+We note that in all three cases, when the target states are located near the middle of the two-exciton manifold, the results exhibit broader population distributions around the target.
+In other words, we find a lower degree of selectivity.
+In comparison, when the target states are located in the lower-energy sector of the two-exciton manifold, we observe a higher degree of selectivity.
+Altering the temporal width of the SPDC pump pulse in the middle column resulted in diminished selectivity, likely due to the activation of multiple pathways.
+In contrast, altering the entanglement time in the right column preserved the selectivity for the majority of states.
 Across all plots, we notice a trend: some two-exciton states remain preferentially populated across all parameter regimes.
-This feature have likely emerged from an intricate interplay of exciton transport parameters leading to the absence of specific relaxation channels.
+This feature has likely emerged from an intricate interplay of exciton transport parameters of mediating states.
 Figure 5: The population distribution for the target state $f_{07}$ which is dominantly excited via the intermediate states $e_{07}$ and $e_{09}$ is displayed for variations in the entangled photon parameters (rows) and for two distinct transport times (columns): $t=0\,\text{fs}$ (immediately after the pulse) and $t=50\,\text{fs}$.
@@
 III.1.2 Simulations: differential preparation of two-exciton populations via preselected mediating states and parametric dependence on photonic properties
-Now, we focus on two specific two-exciton states, $f_{07}$ and $f_{83}$, among the full suite of target states previously monitored, and examine the possibility of their excitation via a chosen pair one-exciton states, selected by setting $\omega_{1}=E_{07}$ and $\omega_{2}=E_{09}$.
-The mediating states are delocalized over sites and prone to ultrafast transport.
-The results are displayed in the left column of Fig.
-4, and Fig.
-5, respectively.
-The entanglement time and width of SPDC pump are set at $T_{\mathrm{ent}}=10\,\mathrm{fs}$ and $\tau_{0}=150\,\mathrm{fs}$, for the upper panel (similar to the reference case in Fig.
-3).
-The width is varied to $\tau_{0}=50\,\mathrm{fs}$ in the middle panel, and the entanglement time is varied to $T_{\mathrm{ent}}=40\,\mathrm{fs}$ in the bottom panel.
-The results demonstrate that the desired narrowband excitation is indeed possible, albeit with low final occupancy.
+We focus on two specific two-exciton states among the full suite of target states previously monitored, namely, $f_{07}$ and $f_{83}$.
+These two-exciton states reside in the lower ($E_{f_{07}}=30192.9\,\text{cm}^{-1}$) and the higher ($E_{f_{83}}=31243.2\,\text{cm}^{-1}$) energy regions of the two-exciton manifold, respectively.
+We examine the possibility of exciting these target states via a fixed pair of one-exciton states, selected by setting $\omega_{1}=E_{e_{07}}=15307.1\,\text{cm}^{-1}$ and $\omega_{2}=E_{e_{09}}=15655.9\,\text{cm}^{-1}$.
+These mediating states are delocalized over multiple chromophoric sites and are prone to ultrafast transport.
+These results are displayed in the left columns of Fig.
+5 and Fig.
+6, respectively.
+The entanglement time and temporal width of the SPDC pump are set to $T_{\mathrm{ent}}=10\,\mathrm{fs}$ and $\tau_{0}=150\,\mathrm{fs}$ for the upper panel, which is similar to the reference case in Fig.
+4.
+In the middle panel, the temporal width is changed to $\tau_{0}=50\,\mathrm{fs}$, whereas in the bottom panel, the entanglement time is changed to $T_{\mathrm{ent}}=40\,\mathrm{fs}$.
+The results demonstrate that the desired narrowband excitation is indeed possible, albeit with low final occupancy in some cases.
 The increase in the temporal width results in a significant loss of selectivity as the number of mediating pathways with comparable spectral weights grows.
@@
 The latter, however, will have consequences as the population redistributes over time, as will be shown in the next section.
+We emphasize that both the target states are excited via the same mediating stating one-exciton states.
+Therefore the differential nature of the final population distribution is dependent on the transition dipoles that connects chosen one-exciton states to the two-exciton target states.
+The final population distribution demonstrates that the short entanglement time reaches a saturation, less so for the $f_{07}$ than $f_{83}$.
+The higher lying transitions for the latter imposes stringent requirements on the frequencies of the entangled photon pairs.
 III.2 Transport in the two-exciton manifold
-The created population distributions undergo time evolution governed by the parameters of the two-exciton transport matrix.
-Here, the following questions are worth investigating: What is the nature of the population evolution for a given state distribution?
-How do two initially created two-exciton population distributions undergo redistribution?
-And how does the nature of redistribution kinetics vary with driving field parameters?
-The kinetic equation that governs the dissipative time evolution of exciton populations, i.e., exciton transport is given by
+The two-exciton population distribution evolves in time, governed by the properties of the transport matrix.
+The following questions merit further investigation: How does the initial two-exciton population distribution redistribute itself over time?
+To what extent the two different population distributions differ after undergoing redistribution?
+Finally, to what extent does the redistribution dynamics depend on the properties of the entangled photon source?
+The kinetic equation that governs the dissipative population dynamics of the two-exciton populations is given by
 $\frac{d}{dt}\rho_{ff}(t)=-\sum_{f^{\prime}}K_{ff,f^{\prime}f^{\prime}}\rho_{f^{\prime}f^{\prime}}(t)$ (13)
-where the transport rate matrices, $K_{ff,f^{\prime}f^{\prime}}$ follow the detailed balance condition $K_{f^{\prime}f^{\prime},ff}/K_{ff,f^{\prime}f^{\prime}}=\exp(-\omega_{f^{\prime}f}/(k_{B}T))$, and $\sum_{f}K_{ff,f^{\prime}f^{\prime}}=0$.
-The initial two-exciton population distribution is obtained via the Eq.
-III.1 which is propagated using the expression: $\rho_{ff}(t_{2})=G_{ff,f^{\prime}f^{\prime}}^{(N)}(t_{2}-t_{1})\rho_{f^{\prime}f^{\prime}}(t_{1})$, where the $G_{f^{\prime}f^{\prime},ff}^{(N)}(\tau)$ is given by first term in Eq.
-II.
+where the transport matrix, $K_{ff,f^{\prime}f^{\prime}}$, follow the detailed balance condition $K_{f^{\prime}f^{\prime},ff}/K_{ff,f^{\prime}f^{\prime}}=\exp(-\omega_{f^{\prime}f}/(k_{B}T))$, and $\sum_{f}K_{ff,f^{\prime}f^{\prime}}=0$.
+The solution to this equation yields the transport Green’s function $\mathcal{G}_{ff,f^{\prime}f^{\prime}}(\tau)$.
+The initial two-exciton population distribution, obtained via Eq.
+III.1, is propagated using the following expression: $\rho_{ff}(t_{2})=\mathcal{G}_{ff,f^{\prime}f^{\prime}}(t_{2}-t_{1})\rho_{f^{\prime}f^{\prime}}(t_{1})$.
+The transport Green’s function is given by the first term in Eq.
+II and is discussed further in Appendix A.
 The following four sets of simulations will help us address the questions posed above.
 III.2.1 Simulations: differential evolution of two-exciton populations and parametric dependence on photonic properties
-We track the differential population evolution following the excitation targeted at $f_{07}$ (right column in Fig.
-4) and $f_{83}$ (right column in Fig.
-5), taken at a delay of $t=50\,\text{fs}$, and $t=250\,\text{fs}$, respectively.
+We monitor the nature of differential population evolution after the excitation targeted at $f_{07}$ (right column in Fig.
+5) and $f_{83}$ (right column in Fig.
+6), taken at a delay of $t=50\,\text{fs}$, and $t=250\,\text{fs}$, respectively.
 Each of the three panels, upper, middle and bottom, represents simulation using different photonic parameters.
@@
 We observe that the evolved distributions at short times retained the memory of prepared distribution in Fig.
-4, but gets washed out in Fig.
-5.
+5, but gets washed out in Fig.
+6.
 The effect of the variation of photonic parameters is also less visible in the latter.
@@
 In Fig.
-6, we track the population redistribution by displaying the population snapshots at $t=50,100,250,1000\,\text{fs}$ following the excitation targeted at $f_{07}$ (in the left column) and $f_{83}$ (in the right column).
+7, we track the population redistribution by displaying the population snapshots at $t=50,100,250,1000\,\text{fs}$ following the excitation targeted at $f_{07}$ (in the left column) and $f_{83}$ (in the right column).
 The results demonstrate two distinct features: the population migrates to the lower energy sector of the manifold rather monotonically for $f_{07}$ and by a slower intra-manifold cascading mechanism for $f_{83}$.
@@
 For detailed discussion, refer to Section III.2
-III.3 Multipoint photon correlation spectroscopy
-The quantity of interest in photon correlation spectroscopy is the time-frequency filtered two-photon coincidence counting signal.
-In this section, we describe how this quantity can monitor the time-evolved population distributions that serve as the initial configuration for the photon emission pathways.
-The signal can be expressed as:
-$S(\{\bar{t}_{j},\bar{\omega}_{j}\})=\int dt_{1}^{\prime}\int d\tau_{1}D^{(1)}(t_{1},\omega_{1};t_{1}^{\prime},\tau_{1})$
+III.3 Photon correlation spectroscopy
+The quantity of interest in photon correlation spectroscopy is the time–frequency filtered two-photon counting signal.
+In this section, we describe how it can monitor the time-evolved population distributions that serve as the initial configuration for the photon emission pathways.
+The rationale for deploying a filtering scheme can be developed as follows.
+Population transport in the two-exciton manifold acts to redistribute the initially created population.
+The timescale of this phenomenon is determined by the duration of transport.
+The population redistribution has two principal consequences: it gives rise to new set of inter-manifold pathways and reduces the spectral intensity of some transitions.
+Each photon emission pathway can be associated with a specific time and frequency window, which can be selectively monitored by tuning the properties of the first set of time and frequency filters.
+Following photon emission, the dynamics continue in the one-exciton manifold, where transport may scramble the population distribution.
+After a finite duration, another inter-manifold transition leads to the ground state.
+Similar to the above, these transitions can also be monitored via appropriate filtering.
+Three pathways in Fig.
+1 (B) show the dynamical complexities accompany each photon emission.
+In systems with exciton manifolds, multiple photon emission pathways are possible.
+The signal is expressed as:
+$S(\bar{t}_{2},\bar{\omega}_{2};\bar{t}_{1},\bar{\omega}_{1})=\int dt_{1}^{\prime}\int d\tau_{1}D^{(1)}(t_{1},\omega_{1};t_{1}^{\prime},\tau_{1})$
 $\times\int dt_{2}^{\prime}\int d\tau_{2}D^{(2)}(t_{2},\omega_{2};t_{2}^{\prime},\tau_{2})\langle N_{r}(t_{1}^{\prime},\tau_{1})N_{s}(t_{2}^{\prime},\tau_{2}^{\prime})\rangle$
@@
 $\sum_{s,s^{\prime}}\sum_{i,i^{\prime}}\langle E_{s^{\prime}}^{\dagger}(t^{\prime}_{2}+\tau_{2})E_{i^{\prime}}^{\dagger}(t^{\prime}_{1}+\tau_{1})E_{i}(t^{\prime}_{1})E_{s}(t^{\prime}_{2})\rangle$ (14)
-where $N(t^{\prime},\tau)=\sum_{s,s^{\prime}}\hat{E}_{sR}^{\dagger}(t^{\prime}+\tau)\hat{E}_{s^{\prime}L}(t^{\prime})$ is a bare photon number operator and $E_{s}(t)=\sqrt{2\pi\hbar\omega_{s}/\Omega}\hat{a}_{s}e^{-i\omega_{s}t}$, and $\Omega$ is the mode quantization volume corresponding to the detector modes.
-We also defined two detector spectrogram functions: $D^{(j)}(\bar{t}_{j},\bar{\omega}_{j};t_{j}^{\prime},\tau_{j})$ that are governed by the parameters $\bar{t}_{j}$ and $\bar{\omega}_{j}$, describing the detector centering times for the temporal, and spectral filter, respectively.
-It can be expressed in terms of the filtering functions as,
+Here, $N(t^{\prime},\tau)=\sum_{s,s^{\prime}}E_{sR}^{\dagger}(t^{\prime}+\tau)E_{s^{\prime}L}(t^{\prime})$ is a photon number operator without the filter.
+The field operators can be expanded as: $E_{s}(t)=\sqrt{2\pi\omega_{s}/\Omega_{s}}\hat{a}_{s}e^{-i\omega_{s}t}$, where $\Omega$ is the mode quantization volume for the photonic modes of the detector.
+The latter follows the Bosonic commutation relation $a_{s},a_{s^{\prime}}^{\dagger}]=\delta_{ss^{\prime}}$.
+We define two spectrogram functions, $D^{(j)}(\bar{t}_{j},\bar{\omega}_{j};t_{j}^{\prime},\tau_{j})$, which are governed by the time and frequency filtering functions, $F_{\bar{t}}^{(j)}$ and $F_{\bar{\omega}}^{(j)}$.
+The spectrogram function can be represented as
 $D(\bar{t},\bar{\omega},t^{\prime},\tau)$
 $=\int\frac{d\omega^{\prime\prime}}{2\pi}e^{-i\omega^{\prime\prime}\tau}|F_{\bar{\omega}}(\omega^{\prime\prime},\omega)|^{2}F_{\bar{t}}^{*}(t^{\prime}+\tau,t)F_{\bar{t}}(t^{\prime},t).$ (15)
-Modulation of these functions individually allows for the examination of the photon emission pathways and the systematic recovery of the photon correlation functions.
-For this work, we choose the Lorentzian filtering functions (see Appendix D), which leads to the spectrogram being expressed as:
+We choose the following time and frequency filters: $F_{\bar{t}}(t^{\prime})=\theta(t^{\prime}-\bar{t})\exp({-\sigma_{\bar{t}}(t^{\prime}-\bar{t})})$ and $F_{\bar{\omega}}(\omega^{\prime})=i(\omega^{\prime}-\bar{\omega}+i\sigma_{\omega})^{-1}$.
+The spectrogram function can be evaluated as
 $D(\bar{t},\bar{\omega},t^{\prime},\tau)=\frac{1}{2\sigma_{\omega}}\theta(t^{\prime}-t)\theta(t^{\prime}+\tau-t)$
 $[\theta(\tau)e^{-\sigma_{\bar{\omega}}\tau}+\theta(-\tau)e^{\sigma_{\bar{\omega}}\tau}]e^{-(i\omega+\sigma_{\bar{t}})\tau-2\sigma_{\bar{t}}(t^{\prime}-t)}$ (16)
-We lay out the filtering strategy: the spectral filters allow state-resolution for the dynamics occurring within similar transport windows; the temporal filters help separate the transport processes that involve transitions with similar energy gaps; combined together, they allow for the tracking of individual state transitions.
-The filtering properties are controlled by the width of the temporal (spectral) filters, $\sigma_{\bar{\omega}_{j}}(\sigma_{\bar{t}_{j}})$, centering times of the temporal (spectral) gate, $\bar{t}_{j}$ ($\bar{\omega}_{j}$).
-The final signal can be expressed as the sum of four pathways (which is reduced to two following symmetry arguments):
-$S(\{\bar{t}_{j},\bar{\omega}_{j}\})$ $=\sum_{p=1}^{4}S^{(p)}(\{\bar{t}_{j},\bar{\omega}_{j}\})=2\text{Re}\sum_{p=1}^{2}S^{(p)}(\{t_{j},\omega_{j}\})$
-$=\sum_{g,e,e^{\prime},f,f^{\prime}}|\mathcal{D}(\omega_{1})\mathcal{D}(\omega_{2})|^{2}|d_{eg}|^{2}|d_{ef}|^{2}\times\int dt_{1}^{\prime}\int d\tau_{1}D_{>}^{(1)}(\bar{t}_{1},\bar{\omega}_{1};t_{1}^{\prime},\tau_{1})\times\int dt_{2}^{\prime}\int d\tau_{2}D_{>}^{(2)}(\bar{t}_{2},\bar{\omega}_{2};t_{2}^{\prime},\tau_{2})$
+It depends on four parameters: the centering times and frequencies of the time and frequency filters, denoted $\bar{t}_{j}$ and $\bar{\omega}_{j}$, and their respective widths, $\sigma_{\bar{t}_{j}}$ and $\sigma_{\bar{\omega}_{j}}$.
+One can modulate the spectrogram by varying these parameters.
+The spectral filters resolve inter-manifold transitions that originate from exciton transport occurring within similar time windows.
+The temporal filters help separate the transport processes that may involve similar spectral gaps.
+Combined, they monitor the inter-manifold transitions jointly in time and frequency space.
+Using the expression of the spectrogram and following the steps described in Appendix E, we can express the signal as a sum of four pathways (which is reduced to two following symmetry arguments)
+$S(\bar{t}_{2},\bar{\omega}_{2};\bar{t}_{1},\bar{\omega}_{1})=\sum_{p=1}^{4}S^{(p)}(\bar{t}_{2},\bar{\omega}_{2};\bar{t}_{1},\bar{\omega}_{1})=2\text{Re}\sum_{p=1}^{2}S^{(p)}(\bar{t}_{2},\bar{\omega}_{2};\bar{t}_{1},\bar{\omega}_{1})$
+$=\sum_{g,e,e^{\prime},f,f^{\prime}}|\mathcal{D}(\bar{\omega}_{1})\mathcal{D}(\bar{\omega}_{2})|^{2}|d_{eg}|^{2}|d_{ef}|^{2}\times\int dt_{1}^{\prime}\int d\tau_{1}D_{>}^{(1)}(\bar{t}_{1},\bar{\omega}_{1};t_{1}^{\prime},\tau_{1})\times\int dt_{2}^{\prime}\int d\tau_{2}D_{>}^{(2)}(\bar{t}_{2},\bar{\omega}_{2};t_{2}^{\prime},\tau_{2})$
 $\langle\langle I|\mathcal{G}_{ge^{\prime},ge^{\prime}}(\tau_{1})\mathcal{G}_{e^{\prime}e^{\prime},ee}(t_{1}^{\prime}-t_{2}^{\prime}-\tau_{2})\mathcal{G}_{ef^{\prime},ef^{\prime}}(\tau_{2})\mathcal{G}_{f^{\prime}f^{\prime},ff}(t_{2}^{\prime})|\rho_{ff}\rangle\rangle$
-$+\sum_{g,e,e^{\prime},f,f^{\prime}}|\mathcal{D}(\omega_{1})\mathcal{D}(\omega_{2})|^{2}|d_{eg}|^{2}|d_{ef}|^{2}\times\int dt_{1}^{\prime}\int d\tau_{1}D_{>}^{(1)}(\bar{t}_{1},\bar{\omega}_{1};t_{1}^{\prime},\tau_{1})\times\int dt_{2}^{\prime}\int d\tau_{2}D_{<}^{(2)}(\bar{t}_{2},\bar{\omega}_{2};t_{2}^{\prime},\tau_{2})$
+$+\sum_{g,e,e^{\prime},f,f^{\prime}}|\mathcal{D}(\bar{\omega}_{1})\mathcal{D}(\bar{\omega}_{2})|^{2}|d_{eg}|^{2}|d_{ef}|^{2}\times\int dt_{1}^{\prime}\int d\tau_{1}D_{>}^{(1)}(\bar{t}_{1},\bar{\omega}_{1};t_{1}^{\prime},\tau_{1})\times\int dt_{2}^{\prime}\int d\tau_{2}D_{<}^{(2)}(\bar{t}_{2},\bar{\omega}_{2};t_{2}^{\prime},\tau_{2})$
 $\langle\langle I|\mathcal{G}_{ge^{\prime},ge^{\prime}}(\tau_{1})\mathcal{G}_{e^{\prime}e^{\prime},ee}(t_{1}^{\prime}-t_{2}^{\prime})\mathcal{G}_{f^{\prime}e,f^{\prime}e}(-\tau_{2})\mathcal{G}_{f^{\prime}f^{\prime},ff}(t_{2}^{\prime}+\tau_{2})|\rho_{ff}\rangle\rangle$ (17)
-The terms $D_{>}$ and $D_{<}$ indicate the sign of the delay $\tau$ variables, and $\mathcal{D}(\omega_{1})\mathcal{D}(\omega_{2})$ is the density of states for the detector, taken at an idealized, narrowband limit around each frequency $\omega_{j}$.
-In our simulation, the influence of this delay was accounted for by evaluating the action of the transport Green’s functions on the initial states within the respective manifolds.
-The rationale for deploying such a scheme can be developed in the following manner.
-During population transport, the ensuing redistribution erases the signatures of the initially created two-exciton wavepacket.
-The timescales of this phenomena depend on the properties of the transport Green’s functions.
-The redistribution give rise to new inter-manifold dipole-allowed transitions and render some transitions diminished in spectral intensity.
-Figure 7: The time-frequency filtered two-photon coincidence counting signal is displayed, showing variations in the widths of both filtering functions and the centering time of the temporal filtering function.
-To control the signal resolution, the following systematic variations relative to the reference simulation (upper-left) are presented: the width of the spectral filter (upper-right); the centering time of the temporal filtering function (middle-left); both the spectral filter width and the temporal filter centering time (middle-right); the width of the temporal filter (bottom-left); and the centering time of the temporal filtering function (bottom-right).
-The targeted excitation of $f_{07}$ by the entangled photon pairs serves as the initial configuration, which is followed by a subsequent series of events: two-exciton manifold transport, inter-manifold transitions, one-exciton manifold transport, and subsequent inter-manifold transitions.
-For a diagrammatic illustration of the processes involved, see Fig.
-2, and for detailed discussion, refer to Section III.3.1
-The subsequent photon emission events can be monitored by appropriately selecting the spectral and temporal filtering properties of the first set of detectors.
-Following the photon emission, the dynamics continue in the one-exciton manifold where transport population scrambling takes place for a time interval, followed by another photon emission event to the ground state.
-Three representative pathways in Fig.1 shows complexities that may be present.
-The joint spectral-temporal filtering strategy can be a powerful tool that enhances the resolution of the bare photon coincidence counting signal.
-Two sets of spectral-temporal filters, defined by the spectrogram can accomplish two distinct tasks: first, the photon emissions events can be temporally windowed to focus on transport timescales and spectrally windowed to focus on the relevant exciton states; second, the photon emissions events can be spectrally windowed such that pairs of inter-manifold transitions originating from the similar two-exciton transport events can be monitored.
-Monitoring these events allows us to construct a kinetic description of dissipative exciton dynamics events triggered by the entangled photon excitation.
-It also help infer how long the remnants of the initially created wavepacket persist.
-The diagrams presented in Figure 2 clarify the physical processes in a more transparent manner.
+In the above, subscripts in the expression for the spectrograms, $D_{>}$ and $D_{<}$, indicate the sign of the delay $\tau$ between two emissions.
+The density of photonic states of the detector is assumed to be concentrated at frequencies corresponding to the centering frequency of the frequency filter; these are denoted as $\mathcal{D}(\bar{\omega}_{1})\mathcal{D}(\bar{\omega}_{2})$.
+In Fig.
+3 (I)–(IV), we diagrammatically illustrate the emission pathways in Liouville space.
+We note that each diagram corresponds to pathways that involve transport in both manifolds, duration of which is controlled by the centering times of the time filters.
+Figure 8: The time-frequency filtered two-photon counting signal is displayed.
+The initial initial population distribution for the photon counting is given by the targeted excitation of $f_{07}$ followed by two-exciton transport.
+Two sets of time and frequency filters can allow monitoring of the subsequent processes such as inter-manifold transition, one-exciton transport, and inter-manifold transition.
+The reference simulation in the upper-left uses the following parameters: width of the frequency filters $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=10\,\text{cm}^{-1}$, width of the time filters $\sigma_{\bar{t}_{1}}=\sigma_{\bar{t}_{2}}=4.8681\,\text{cm}^{-1}$, waiting times $t_{\text{w},1}=0\,\text{fs}$, and $t_{\text{w},2}=100\,\text{fs}$.
+In upper-right the the width of the frequency filter is varied to $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$.
+In middle-left the centering time of the time filter is varied to capture $t_{w,2}=1000\text{fs}$, whereas in middle-right we vary widths of frequency filters $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$ while keeping the and the centering time of the time filter same as middle-left; In bottom-left, the width of the time-filter is varied to $\sigma_{\bar{t}_{2}}=0.5409\,\text{cm}^{-1}$.
+In bottom-right, the centering time of the time filter is varied to capture $t_{w,1}=50\text{fs}$.
+For detailed discussion, see Section III.3.1
 III.3.1 Simulation: two-photon coincidence counting signal with variation in filtering parameter and different waiting times
 In Fig.
-7, we present six two-dimensional plots illustrating the signal where the filtering widths and centering parameters (the latter relates to the transport times) are varied.
-The centering times of the spectral filters are taken as the scanning parameters.
-In these plots, a series of features can be noticed: the peaks along the $\bar{\omega}_{1}$ axis correspond to transitions from the two to and one-exciton manifold ($\sim\omega_{fe}$), while the peaks along the $\bar{\omega}_{2}$ axis correspond to transitions from the one-exciton manifold to the ground state ($\sim\omega_{eg}$).
-The one-exciton manifold has a spectral width of $\approx 950\,\text{cm}^{-1}$.
-The density of energy gaps is significantly higher in the upper set of transitions.
-Even though, for the chosen example, the initial state distribution, following excitation of $f_{07}$, effectively spans only the first ten levels of the two-exciton manifold (an energy range of $\approx 350\,\text{cm}^{-1}$), the number of possible $\omega_{fe}$ transitions remains high.
-The upper-left panel (the reference simulation for this set) employs the following filtering parameters: spectral widths $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=10\,\text{cm}^{-1}$, temporal widths $\sigma_{\bar{t}_{1}}=\sigma_{\bar{t}_{2}}=4.8681\,\text{cm}^{-1}$, and two-exciton (one-exciton) waiting times $t_{\text{w},1}=0\,\text{fs}$ ($t_{\text{w},2}=100\,\text{fs}$).
-Variables $t_{\text{w},j}$ are associated with the $t_{2},t_{1}$ integrals when the detector spectrograms are inserted into the relevant expression.
-We observe eight resonance zones spread along the $\bar{\omega}_{1}(/10^{3}\,\text{cm}^{-1})$ axis, each containing several resonances that can be resolved along the $\bar{\omega}_{2}(/10^{3}\,\text{cm}^{-1})$ axis.
-For easier interpretation, we may classify these resonances along $\bar{\omega}_{1}$ as four distinct clusters, containing the following number of resonance zones: two ($\sim 14.0-14.2$), three ($\sim 14.3-14.7$), two ($\sim 14.9-15.1$), and one ($\sim 15.3-15.5$).
-Each of these clusters may have multiple origins in the upper manifold.
-The four observed $\bar{\omega}_{1}$ resonance clusters correspond approximately to the transitions $\omega_{f_{1}e_{13}}$ and $\omega_{f_{2}e_{10}}-\omega_{f_{2}e_{13}}$; $\omega_{f_{1}e_{8}},\omega_{f_{2}e_{6}},\omega_{f_{2}e_{5}}$; $\omega_{f_{4}e_{3}}$ and either $\omega_{f_{3}e_{1}}$ or $\omega_{f_{4}e_{1}}$; and $\omega_{f_{7}e_{1}}-\omega_{f_{8}e_{1}}$, respectively.
-Taking each cluster, the $\bar{\omega}_{2}$ ($10^{3}\,\text{cm}^{-1}$) resonances are located at $\sim 15.5,15.8$ and $\sim 15.4,15.7$ (corresponding to $\omega_{e_{8}g},\omega_{e_{12}g}$ and $\omega_{e_{8}g},\omega_{e_{12}g}$); $\sim 15.3,15.6$, $\sim 15.3,15.4$, and $\sim 15.3,15.4$ ($\omega_{e_{7}g},\omega_{e_{8}g}$; $\omega_{e_{7}g},\omega_{e_{8}g}$; and $\omega_{e_{7}g},\omega_{e_{8}g}$); $\sim 15.1,15.3$ ($\omega_{e_{1}g},\omega_{e_{3}g}$); and $\sim 15.0$ ($\omega_{e_{1}g}$), respectively.
-This plot will serve as a reference simulation.
-The upper-right panel presents the results of varying the spectral filter width by setting $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$, while keeping all other parameters fixed to those of the reference simulation.
-We observe that several resonance zones in the second cluster are absent along the $\bar{\omega}_{1}$ axis, with the most noticeable absences occurring in the middle and lower energy sectors.
-Additionally, among the present zones, a few resonances along the $\bar{\omega}_{2}$ axis are absent, and the fourth cluster has been diminished in intensity.
-This observation may be rationalized by noting the increased destructive interferences permitted by the broadband spectral filter.
-The middle-left panel investigates the effect of longer waiting times, specifically using $t_{\text{w},2}=1000\,\text{fs}$ for the one-exciton transport, while keeping all other parameters unchanged from the reference simulation.
-The population distribution in the one-exciton manifold is dependent on the cascaded emission from the two-exciton manifold.
-We note that the clusters have further coalesced along the $\bar{\omega}_{1}$ axis and are clearly separated into two dominant regions.
-The latter is indicative of the dipole transition rule dominating the spectral weight at these longer times.
+8, we present six two-dimensional plots where centering times of the frequency filters (i.e., $\bar{\omega}_{j}$) are taken as the scanning parameters.
+These results explore the signal for cases where the widths (i.e., $\sigma_{\bar{t}_{j}}$) and centering parameters of time filters (i.e., $\bar{t}_{j}$) are varied.
+We note that, the latter help unravel the exciton transport timescales.
+The two-exciton distribution targeted at $f_{07}$, after a period of time evolution serves as the initial configuration.
+The peaks along the $\bar{\omega}_{1}$ axis correspond to transitions from the two-exciton to and one-exciton manifold $|f\rangle\rightarrow|e\rangle$, while the peaks along the $\bar{\omega}_{2}$ axis correspond to transitions from the one-exciton manifold to the ground state $|e\rangle\rightarrow|g\rangle$.
+The former monitors the spectral gaps $\sim\omega_{fe}$ while the latter the spectral gaps $\sim\omega_{eg}$.
+This can be ensured by choosing the centering times of the time filter $\bar{t}_{2}$ well-ahead of $\bar{t}_{1}$.
+As highlighted earlier in Section II, density of energy gaps is significantly higher in the upper set of transitions.
+We recall that the one-exciton manifold hosts $14$ states within a spectral width of $\approx 970\,\text{cm}^{-1}$ and following excitation of $f_{07}$, the two-exciton manifold hosts $\sim 10$ states within a spectral width of $\approx 350\,\text{cm}^{-1}$.
+The number of plausible transitions remains large $\sim 140$.
+Two filtering stages allow classify these transitions in different temporal windows.
+For the upper-left panel, the reference simulation for this set, we use the following parameters: $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=10\,\text{cm}^{-1}$, $\sigma_{\bar{t}_{1}}=\sigma_{\bar{t}_{2}}=4.8681\,\text{cm}^{-1}$, waiting times $t_{\text{w},1}=0\,\text{fs}$, and $t_{\text{w},2}=100\,\text{fs}$.
+Variables $t_{\text{w},j}$ are associated with the $\tau_{j}$ integrals in the expression of detector spectrograms.
+The waiting times determine the duration for which exciton transport is allowed at each manifold, before the time-filter is opened.
+In the plot, we observe eight resonance zones spread along the $\bar{\omega}_{1}$ axis, each containing several resonances that can be further resolved along the $\bar{\omega}_{2}$ axis.
+For easier interpretation, we may classify the resonances along $\bar{\omega}_{1}/10^{3}\,\text{cm}^{-1}$ in four distinct clusters, each containing two ($\sim 14.0-14.2$), three ($\sim 14.3-14.7$), two ($\sim 14.9-15.1$), and one ($\sim 15.3-15.5$) resonances.
+Since only the transition energy differences are being captured, each of these transitions, in the absence of the time filter would have appeared in a congested form.
+The four observed clusters correspond approximately to the transitions $\omega_{f_{1}e_{13}}$ and $\omega_{f_{2}e_{10}}-\omega_{f_{2}e_{13}}$; $\omega_{f_{1}e_{8}},\omega_{f_{2}e_{6}},\omega_{f_{2}e_{5}}$; $\omega_{f_{4}e_{3}}$ and either $\omega_{f_{3}e_{1}}$ or $\omega_{f_{4}e_{1}}$; and $\omega_{f_{7}e_{1}}-\omega_{f_{8}e_{1}}$, respectively.
+For each of these clusters, the resonances along $\bar{\omega}_{2}/10^{3}\,\text{cm}^{-1}$ are identified as $\sim 15.5,15.8$ and $\sim 15.4,15.7$ (corresponding to $\omega_{e_{8}g},\omega_{e_{12}g}$ and $\omega_{e_{8}g},\omega_{e_{12}g}$); $\sim 15.3,15.6$, $\sim 15.3,15.4$, and $\sim 15.3,15.4$ ($\omega_{e_{7}g},\omega_{e_{8}g}$; $\omega_{e_{7}g},\omega_{e_{8}g}$; and $\omega_{e_{7}g},\omega_{e_{8}g}$); $\sim 15.1,15.3$ ($\omega_{e_{1}g},\omega_{e_{3}g}$); and $\sim 15.0$ ($\omega_{e_{1}g}$), respectively.
+The upper-right panel repeats the simulation with $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$, while keeping all other parameters fixed to those of the previous one.
+We observe, along the $\bar{\omega}_{1}$ axis, that several resonance zones in the second cluster are absent; the most noticeable absences occur in the middle and low energy sectors.
+Additionally, a few resonances along the $\bar{\omega}_{2}$ are absent.
+The fourth cluster has been diminished in intensity.
+The broadband spectral filters used in the simulation increases the number of destructive interferences.
+Given that the transitions originate from a limited number of states number of states it also highlights that they are preferentially transitioning to a few selected one-exciton states.
+The middle-left panel investigates the effect of longer waiting times aimed at identifying the role of one-exciton transport.
+We set $t_{\text{w},2}=1000\,\text{fs}$ while keeping all other parameters unchanged from the reference simulation.
+The population distribution in the one-exciton manifold depends on two factors: the dipole-connectivity of states with those of the two-exciton manifold and the transport rates.
+We note that along the $\bar{\omega}_{1}$ axis the clusters have further coalesced.
+They are clearly separated into two dominant regions.
 Furthermore, the fourth cluster has vanished, and the resolution along the $\bar{\omega}_{2}$ axis is lost.
-The middle-right panel explores the variation of resolution compared to the middle-left one by varying both the spectral filter width and the one-exciton waiting time ($t_{\text{w},2}$).
-The parameters are set jointly equal to those used in the two previous simulations ($\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$ and $t_{\text{w},2}=1000\,\text{fs}$), while keeping all other parameters fixed.
-We observe that widening the spectral gate may not have permitted a larger number of resonance structures along the $\bar{\omega}_{2}$ axis.
-This implies that a broader set of transitions remain unaccounted for (either not presented or not dipole allowed).
-Importantly, these transitions, despite differing widely in energy, participate in similar dynamics due to the longer transport window afforded by the extended waiting time.
-The lower panels are dedicated to exploring two distinct cases: the bottom-left panel examines the role of the width of the temporal filter (by setting $\sigma_{\bar{t}_{2}}=0.5409\,\text{cm}^{-1}$ ), and the bottom-right panel examines the role of the two-exciton transport waiting time (by setting $t_{\text{w},1}=50\,\text{fs}$), while keeping other parameters the same as the reference.
-For the bottom-left panel, we note that the second cluster has a diminished presence, with several resonances along the $\bar{\omega}_{2}$ axis being absent, an effect attributed to the shorter temporal filter effectively skipping some transitions occurring at longer respective timescales.
-For the bottom-right panel, in contrast with the reference simulation, we observe that the double resonance structures along the $\bar{\omega}_{2}$ axis have coalesced into single structures, while there is no significant change along the $\bar{\omega}_{1}$ axis.
-This can be attributed to the fact that at longer two-exciton transport duration, state-specific transport rates exhibited minor differences.
-The redistributed the spectral weights, in turn, has resulted in a fewer $\omega_{fe}$ transitions becoming dominant.
-The initial states in all cases can be tracked from Figure 6. However, the state indices require a conversion to the energy domain (energy gaps) for constructing a robust phenomenological description.
+The observation indicates that, at longer waiting times, the connectivity dictated by transition dipoles focuses the spectral weights to a few selected states.
+The middle-right panel repeats the reference simulation but with $\sigma_{\bar{\omega}_{1}}=\sigma_{\bar{\omega}_{2}}=20\,\text{cm}^{-1}$ and $t_{\text{w},2}=1000\,\text{fs}$.
+Therefore it is aimed at exploring the role of a broadband frequency filter on one-exciton transport at longer times.
+We note the resonances along the $\bar{\omega}_{2}$ axis and conclude that widening the spectral gate may not have permitted a larger number of transitions.
+The reasons are twofold: the initial two-exciton distributions, being restricted to a limited set of states, have no further transitions to explore and these transitions at longer times participates in dynamics that converges to same set of states.
+The bottom-left panel explores the examines the role of the width of the time filter by setting $\sigma_{\bar{t}_{2}}=0.5409\,\text{cm}^{-1}$ while keeping all other parameters the same as the reference.
+We note that the second cluster has a diminished presence.
+Also, several resonances along the $\bar{\omega}_{2}$ axis are absent.
+We note that positioning the centering time of the filter at shorter time effectively skips the resonances that are generated as a result of the longer time dynamics.
+Hence, by comparing the reference simulation and this one, one may identify the transitions that takes place at longer times.
+The bottom-right panel examines the role of the two-exciton transport by setting $t_{\text{w},1}=50\,\text{fs}$.
+All other parameters are kept as the same as that of the reference.
+We observe that the double resonance structures along the $\bar{\omega}_{2}$ axis have coalesced into single structure, while there is no significant change along the $\bar{\omega}_{1}$ axis.
+The longer two-exciton transport tend to funnel all populations to a few low-lying two-exciton states, as can be observed in Fig.
+5 and Fig.
+7.
+This makes a fewer number of transitions possible, reducing the number of corresponding resonances.
+We note that the for a robust phenomenological description the state indices from the figures need to be converted to energy values.
 IV Conclusions
-In this article, we had demonstrated that a combined protocol utilizing entangled photon-pair excitation and temporally-spectrally filtered photon counting provides a systematic route to probe dissipative multi-exciton kinetics.
-The protocol, illustrated via simulations for the exciton aggregate LHCII, is presented across three subsequent sections.
-It revealed, in section III.1, that narrow-band distribution around a special class of two exciton states can indeed be excited with considerable degree of control.
-The final population distribution is benchmarked against the variation of the entangled photon parameters in section III.1.2.
-Here we examined the pathway selectivity by targeting two different two exciton state that are excited via two pre-selected one-exciton states.
-The exciton-phonon interactions in multi-site molecular aggregates are characterized by exciton transport and dephasing dynamics spanning multiple time and energy scales.
-The entangled photon pairs demonstrated a superior ability maintain pathway selectivity.
-Following the excitation, and population evolution, illustrated in section III.2.1-III.2.2), we introduced the detection scheme to monitor the fate of the time-evolving population, and classify the emission pathways.
-It was shown, in section III.3, that the joint spectral and temporal filtering of the emitted fluorescence allow a more nuanced monitoring of inter-manifold transitions competing with intra-manifold transport kinetics.
-We note that the targeted states, in III.1.2, are chosen due to their higher overlap with the local overtone states.
-These properties can be quantified by analyzing the transformation matrix elements, $T^{(2)}_{k,mm}$.
-While state-selective excitations in interacting molecular aggregates conventionally yield spatially delocalized wavepackets, an important corollary suggests that a specially crafted exciton wavepacket, that is, a tailored superposition of exciton eigenstates, may allow for the excitation of spatially localized wavepackets.
-Such localization could then enhance the propensity of exciton-exciton annihilation provided that significant non-adiabatic coupling elements between the higher-lying states allow the internal conversion to occur.
+In this work, we demonstrated that a combining entangled photon-pair excitation and time-frequency filtered two-photon counting provides a systematic route to probe dissipative one- and two-exciton kinetics.
+The combined protocol was applied to the interacting molecular aggregate, light-harvesting complex II (LHCII), where exciton–exciton and exciton–phonon interactions jointly govern to the dynamics.
+The protocol offers a classification scheme for exciton dynamics, even when they span multiple time and frequency scales.
+Results in the Section III.1 reveals that a narrow-band population distribution around some of the two-exciton states can be created with a high degree of fidelity.
+The final population distribution is examined against variations in the entangled photon parameters and the temporal width of the SPDC pump.
+We also examined the differential nature of excitation by targeting two distinct two-exciton states.
+Because these two target states are excited via the same one-exciton states, they offer insight into pathway selectivity.
+From the current simulations, a strategy emerges: preferentially accessing the one-exciton resonances that are less susceptible to transport is the key to achieving robust, state-selective excitation of two-exciton states.
 The expression in Eq.
-III.1 remains valid for a wide range of correlated optical fields, albeit in the weak-field regime.
-In the expression, five terms appear, each of which contains a convolution of inter-manifold coherence and intra-manifold population Green’s functions with four-point photon correlation functions.
-The latter can be interpreted as pathway weights, and multiobjective optimization schemes can be used to actively design them.
-Future work will address this issue.
-Similarly, the expressions in Eq.
-III.3 remain valid for any general initial state distributions.
-The variation of the filtering parameters offers us the ability to classify inter-manifold transitions in the presence and competing with intra-manifold transport kinetics.
-It allows better assignment timescales and generate a dynamical picture for one and two-exciton transport kinetics.
-We assumed that the excitation and emission events are temporally separated, omitting spontaneous Raman contributions.
-From these simulations, a probing strategy may emerge: preferentially accessing the one-exciton resonances that are less susceptible to transport within a given temporal window may be the key to achieving controlled excitation.
-We note that the full protocol encountered transport events thrice during its course.
-Within the model and the chosen expressions, the entanglement time, controlled by the length of the birefringent crystal, determines the temporal window of the photon-exciton interactions.
-The bandwidth of the classical SPDC pump, on the other hand, determines the range of accessible resonances.
-The independent controllability of these variables affords some flexibility in selecting their values.
-In turn, it will allow control over a broad range of time and frequency parameter domain.
-Analogously, the choice of filtering parameters, decided independently, must take into account the excitation properties to be able to characterize a wide range of two-photon emission pathways.
-The utility of correlated photonic sources for multiphoton spectroscopies promises two main advantages in the appropriate parameter regime: the ability to actively modulate multiple material pathways to highlight unseen spectral signatures, and favorable scaling of the signal with the incoming field intensity.
-However, these benefits face a significant challenge due to the inherently weak signal-to-noise ratio.
-Technological advances in this direction would make the proposed protocol a suitable investigative tool for exploring potential non-classical effects in light-harvesting complex.
-Alternatively, the scheme remains applicable in the high-gain regime using squeezed photon sources.
+III.1 remains valid for a wide range of photonic sources, provided the dynamics remains within the perturbative regime of exciton–photon coupling.
+The expression consists of five terms, each containing a convolution of exciton Green’s functions with four-point correlation functions of the driving photonic sources.
+They modulate spectral weights for each pathway.
+Our future work will address how multiobjective optimization algorithms can be used to actively tailor these correlation functions.
+The evolution of the initially created two-exciton population is studied in Section III.2.
+This evolved population distribution serves as the initial configuration for the photon emission pathways.
+The presence of transport within the two-exciton manifold effectively scrambles the identity of the states initially created by two-photon absorption.
+Similarly, transport in the one-exciton manifold have the same effects on the state distribution generated after the first photon emission.
+Interestingly, the one-exciton transport may exert the scrambling effect on two occasions: during the excitation and again during the emission process.
+In Section III.3, we deployed the time-frequency filtering to track inter-manifold transitions.
+The expressions in Eq.
+III.3 remain valid for general two-exciton population distributions.
+The variation of the filtering parameters may offer detailed insights into spectral gaps, intra-manifold transport timescales, and inter-manifold dephasing.
+We assumed that the excitation and emission events are temporally separated which omits spontaneous Raman contributions.
+The joint spectral-temporal filtering effectively enhances the resolution of the bare two-photon counting signal.
+Two spectrograms, parametrized by filtering functions separate the exciton pathways jointly in time and frequency.
+It allows one to construct a kinetic description of state-resolved exciton dynamics even in the presence of dissipation.
+The entanglement time, controlled by the length of the birefringent crystal, determines the temporal window of successive photon-exciton interactions.
+In contrast, the bandwidth of the classical spontaneous parametric down-conversion (SPDC) pump determines the range of accessible resonances.
+The independent controllability of these variables may affords additional flexibility.
+For a given $N$-site exciton aggregate, there are $N$ one-exciton states and N+N(N1)/2N+N(N-1)/2 two-exciton states.
+As the size of the aggregate increases, the dynamics may span a broad range of time and frequency values.
+To probe such dynamics a wider tunability of photonic parameters may become necessary [5, 37, 11].
+Similar constraints apply on the filtering parameters, which must be able to characterize a wide range of two-photon emission pathways.
+In general, entangled photon pairs offer two primary advantages for two-photon spectroscopy: the ability to modulate competing excitation (and de-excitation) pathways to selectively suppress or amplify specific exciton resonances, and favorable scaling of the signal with the intensity of the incoming field.
+We note that the correlation function of coherent laser fields factorizes into a product of four amplitude functions, i.e., $\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle=A^{*}_{4}(\omega_{4})A^{*}_{3}(\omega_{3})A_{2}(\omega_{2})A_{1}(\omega_{1})$.
+The signal scales with the square of the incident field intensity.
+In contrast, for the entangled photon pairs, the signal scales linearly.
+However, realizing these benefits remains a significant challenge due to the inherently low signal-to-noise ratio and the difficulty of separating contributions arising from photonic entanglement [45, 26, 27].
+Technological advances in this direction would establish the proposed protocol as a suitable tool for investigating non-classical effects in light-harvesting complexes.
+The time- and frequency-filtered $n$-photon counting signal can analyze multiphoton emission in a $2n$-dimensional parameter space.
+If combined with an efficient state preparation using multiphoton entangled sources, it may provide a robust spectroscopic tool for characterizing dissipative molecular excitonic systems.
+Such approaches can also be extended to programmable photonic sources and interferometric detection schemes [28, 13, 3].
 Acknowledgements.
@@
 The Brownian oscillators are characterized by the Huang-Rhys parameter, mode frequency, and correlation time, denoted as $\lambda_{j}$, $\gamma_{j}$, $\omega_{j}$, respectively.
-Relevant parameter values are detailed in the [9, 8].
-The phonon interaction gives rise to exciton transport (first term) and dephasing (second term) Eq. II. The exciton transport components can be expressed as
-$[\exp{(-K^{(n)}t)}]_{a_{1}a_{1},a_{3}a_{3}}$ $=\sum_{p}\chi_{a_{1}p}^{R}D_{pp}^{-1}e^{-\lambda_{p}t}\chi_{pa_{3}}^{L}$ (21)
-where $\lambda_{p}$ is the p-th eigenvalue of the transport matrix, $\chi^{L}(\chi^{R})$ are the left (right) eigenvectors $K\chi^{R}=\lambda_{p}\chi^{R}$ and $\chi^{L}K=\lambda_{p}\chi^{L}$ of the same with additional relation $D_{pp}=\chi^{L}\chi^{R}$.
-The inter-manifold coherence green’s functions, contains excitonic resonances $i\omega_{a_{1}a_{2}}^{(n)}-\gamma_{a_{1}a_{2}}^{(n)}t=z_{a_{1}a_{2}}^{(n)}$, where the line-broadening function term is evaluated as
-$\gamma_{a_{j}}^{(n)}(t)$ $=\sum_{a_{j^{\prime}}}C_{\mathrm{ph}}(\omega_{a_{j}a_{j^{\prime}}}^{(n)})\sum_{m}T_{ma_{j}}^{(n)}T_{ma_{j}}^{(n)}T_{ma_{j^{\prime}}}^{(n)}T_{ma_{j^{\prime}}}^{(n)}$ (22)
-and the phonon correlation function evaluated at resonant exciton transition frequencies is given by
-$C_{\text{phonon}}(\Omega)=\int_{0}^{\infty}dt\exp(i\Omega t)$
-$\qquad\times\int\frac{d\omega}{2\pi}J(\omega)\Big[\coth{(\beta\omega/2)}\cos{\omega t}\mp i\sin{\omega t}\Big]$ (23)
-Therefore characteristic broadening of the resonances is determined by the phonon spectral function, and exciton gaps.
-Appendix B Note on the observable: two-exciton population distribution
+Relevant parameter values are detailed in the [10, 9, 41, 40].
+The phonon interaction gives rise to exciton transport (first term) and dephasing (second term) Eq. II. The exciton transport Green’s function can be expressed as
+$\mathcal{G}_{a_{1}a_{1},a_{3}a_{3}}(t)$ $=[\exp{(-Kt)}]_{a_{1}a_{1},a_{3}a_{3}}$
+$=\sum_{p}\chi_{a_{1}p}^{R}D_{pp}^{-1}e^{-\lambda_{p}t}\chi_{pa_{3}}^{L}$ (19)
+where $\lambda_{p}$ is the p-th eigenvalue of the transport matrix, $\chi^{L}(\chi^{R})$ are the left (right) eigenvectors that satisfy $K\chi^{R}=\lambda_{p}\chi^{R}$ and $\chi^{L}K=\lambda_{p}\chi^{L}$ of the same with additional relation $D=\chi^{L}\chi^{R}$.
+The intra-manifold exciton coherence Green’s function is given by
+$\mathcal{G}_{a_{1}a_{2}}(t)$ $=\exp{[-i\omega_{a_{1}a_{2}}t-\gamma_{a_{1}a_{2}}t]}$ (20)
+where we defined the pure dephasing parameter
+$\tilde{\gamma}_{a_{1}a_{2}}=\gamma_{a_{1}a_{2}}-\frac{1}{2}(K_{a_{1}a_{1},a_{1}a_{1}}+K_{a_{2}a_{2},a_{2}a_{2}}).$ (21)
+and the dephasing parameters $\gamma_{a_{1}a_{2}}=-K_{a_{1}a_{2},a_{1}a_{2}}$.
+Figure 9: Histograms describing distribution of spectral gaps corresponding for $\omega_{eg}$, $\omega_{fg}$, and $\omega_{fe}$ at the top, middle and bottom panel respectively.
+Each energy bin have a width of $200\,\text{cm}^{-1}$.
+An increase in the number of counts indicates corresponding increase in inter-manifold transitions.
+For a detailed discussion, refer to Appendix B.
+The state-specific line-broadening functions are defined as
+$\gamma_{a_{1}}(t)$ $=\sum_{a_{2}}C(\omega_{a_{1}a_{2}})\sum_{m}T_{ma_{1}}T_{ma_{1}}T_{ma_{2}}T_{ma_{2}}$ (22)
+The dissipation function, $C(\omega_{ab})$, is the Fourier transform of the phonon correlation function evaluated at exciton transition frequencies $\omega_{ab}$, expressed as
+$C(\omega_{ab})=\int_{0}^{\infty}dt\exp(i\omega_{ab}t)$
+$\qquad\times\int\frac{d\omega}{2\pi}J_{\text{ph}}(\omega)\Big[\coth{(\beta\omega/2)}\cos{\omega t}\mp i\sin{\omega t}\Big]$ (23)
+It indicates that the line broadening functions are determined by the phonon spectral function, and exciton gaps.
+The inter-manifold coherence Green’s functions, contains excitonic resonances of the type: $iz_{a_{1}a_{2^{\prime}}}=\omega_{a_{1}a_{2^{\prime}}}-i\gamma_{a_{1}a_{2^{\prime}}}t$, where the primed indices indicate states belonging to different manifolds.
+The values of the inter-manifold dephasing are estimated from the expression of the line broadening by replacing respective manifold terms with primed indices.
+Correlations between the source of driving and exciton-phonon dissipation was neglected [12].
+Appendix B Note on the exciton energy distribution: one and two-exciton manifold
+As stated in section II), the exact-diagonalization step yields $14$, one exciton states (denoted $|e_{j}\rangle$) and $105$ two-states (denoted $|f_{k}\rangle$).
+The distribution of transition energy gaps or spectral gaps for each of these manifolds can be evaluated using a pair of exciton energies belonging to different manifolds.
+For transitions between the ground and one-exciton states, ranging from $\omega_{e_{1}g_{1}}=E_{e_{1}}$ to $\omega_{e_{14}g_{1}}=E_{e_{14}}$, there are $14$ spectral gaps.
+The same between the ground and two-exciton states are $105$, which ranges from $\omega_{f_{1}g_{1}}=E_{f_{1}}$ to $\omega_{f_{105}g_{1}}=E_{f_{14}}$.
+In comparison, spectral gaps between the one and two exciton manifold spans from $\omega_{f_{1}e_{14}}=13909.6\,\text{cm}^{-1}$ to $\omega_{f_{105}e_{1}}=16847.3\,\text{cm}^{-1}$ contributing to a $14\times 105$, a polynomial large number of spectral gaps.
+A majority of them are dipole-allowed.
+These gaps appear in the expressions of exciton resonances via Green’s functions.
+The distribution of spectral gaps are visualized via three histograms in Fig.
+9.
+In the Fig.
+4, the simulation considered a special situation: the pump photon energies are downconverted to degenerate entangled photon pairs.
+The energies of the sequentially corresponds to the each spectral gaps corresponding to each state in the the two-exciton manifold.
+For the chosen two examples of two exciton states $f$ and $f$, described in
+Appendix C Note on the observable: two-exciton population distribution
+Here, we present an outline for the derivation of the signal presented in Section III.1.
 The observable is defined using the projection operator onto the desired two-exciton state, averaged over the time-dependent density operator.
 Assuming the perturbative regime of exciton-driving field interaction to hold, the density operator can be expanded to the fourth-order in incoming field-exciton interaction Hamiltonian to obtain
-$O_{ff}(t)$ $=(-i)^{4}\text{Re}\int^{\infty}_{-\infty}d\tau_{4}\int^{\tau_{4}}_{-\infty}d\tau_{3}\int^{\tau_{3}}_{-\infty}d\tau_{2}\int^{\tau_{4}}_{-\infty}d\tau_{1}\Big\langle P_{+}(t)H_{\text{int},R}(\tau_{3})H_{\text{int},R}(\tau_{2})H_{\text{int},L}(\tau_{4})H_{\text{int},L}(\tau_{1})\Big\rangle$
+$\rho_{ff}(t)$ $=(-i)^{4}\text{Re}\int^{\infty}_{-\infty}d\tau_{4}\int^{\tau_{4}}_{-\infty}d\tau_{3}\int^{\tau_{3}}_{-\infty}d\tau_{2}\int^{\tau_{4}}_{-\infty}d\tau_{1}\Big\langle P_{+}(t)H_{\text{int},R}(\tau_{3})H_{\text{int},R}(\tau_{2})H_{\text{int},L}(\tau_{4})H_{\text{int},L}(\tau_{1})\Big\rangle$
 $=(-i)^{4}\text{Re}\int^{\infty}_{-\infty}dt_{4}\int^{\infty}_{0}dt_{3}\int^{\infty}_{0}dt_{2}\int^{\infty}_{0}dt_{1}\langle P_{+}\mathcal{G}(t_{4})d_{R}\mathcal{G}(t_{3})d_{R}\mathcal{G}(t_{2})d_{L}^{\dagger}\mathcal{G}(t_{1})d^{\dagger}_{L}\rangle$
 $\qquad\langle E^{\dagger}(t-t_{4})E^{\dagger}(t-t_{4}-t_{3})E(t-t_{4}-t_{3}-t_{2})E(t-t_{4}-t_{3}-t_{2}-t_{1})\rangle$ (24)
-In the expression using superoperator notation [37, 12], the variables $\tau_{j}$ denoted the exciton-photon interaction times in the first line.
-It is converted into time-delays $t_{j}=\tau_{ij}=\tau_{i}-\tau_{j}$ in the subsequent line.
-The field expression correlate the exciton evolution via the time integrals.
-Expanding the exciton correlation function in the basis of one and two-exciton states, and carrying out the implicit averaging over the phonons lead to an computable expression of the signal which contains the exciton Green’s functions (defined in Eq.
-II).
-We further recast the expression in the frequency domain to obtain
-$O_{ff}(t)$ $=2\,\text{Re}\Big\{\int\frac{d\omega_{4}}{2\pi}\int\frac{d\omega_{3}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}}{2\pi}e^{+i\omega_{3}\tau_{3}+i\omega_{4}\tau_{4}-i\omega_{2}\tau_{2}-i\omega_{1}\tau_{1}}\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
+In the expression, the variables $\tau_{j}$ and $t_{j}=\tau_{ij}=\tau_{i}-\tau_{j}$ denote the exciton-photon interaction times and time-delays, respectively.
+The four-point correlation function of the driving field correlate the four-point correlation function of the exciton operators via the time integrals.
+Expanding the exciton correlation function in the basis of one and two-exciton states, carrying out the averaging over the phonons and inserting the frequency domain variables lead the expression of the signal
+$\rho_{ff}(t)$ $=2\,\text{Re}\Big\{\int\frac{d\omega_{4}}{2\pi}\int\frac{d\omega_{3}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}}{2\pi}e^{+i\omega_{3}\tau_{3}+i\omega_{4}\tau_{4}-i\omega_{2}\tau_{2}-i\omega_{1}\tau_{1}}\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
 $\Big\{d_{fe^{\prime}}d_{e^{\prime}g}d_{fe}^{*}d_{eg}^{*}\times\mathcal{G}_{ff}(-\omega_{4}-\omega_{3}+\omega_{2}+\omega_{1})\mathcal{G}_{fe^{\prime}}(-\omega_{3}+\omega_{2}+\omega_{1})\mathcal{G}_{fg}(\omega_{2}+\omega_{1})\mathcal{G}_{eg}(+\omega_{1})$
@@
 $+d_{fe}^{*}d_{fe^{\prime}}d_{e^{\prime}g}d_{eg}^{*}\times\mathcal{G}_{ff}(+\omega_{2}-\omega_{4}-\omega_{3}+\omega_{1})\mathcal{G}_{ef}(-\omega_{4}-\omega_{3}+\omega_{1})\mathcal{G}_{e^{\prime}e}(-\omega_{3}+\omega_{1})\mathcal{G}_{eg}(+\omega_{1})\Big\}\Big\}$ (25)
-In expression above is significantly more detailed.
-Here five types of phonon-averaged exciton Green’s functions, representing one, and two-exciton inter-manifold coherence, one and two-exciton transport and one-exciton intra-manifold coherence appear.
-They also bear accumulated frequencies of the driving field in their arguments.
+The expression contains three types of exciton Green’s functions that describes inter-manifold coherence, intra-manifold coherence and transport.
 The first term represents the two-exciton coherence-mediated pathways, the second and fourth terms represents one-exciton transport-mediated pathways and the third and fifth terms represents the one-exciton coherence-mediated pathways.
-A more explicit form of the expression can be obtained if we plug in the expressions of the exciton Green’s functions.
-It gives
-$O_{ff}(t)$ $\propto 2\mathrm{Re}\Big\{\int\frac{d\omega_{4}}{2\pi}\int\frac{d\omega_{3}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}}{2\pi}e^{+i\omega_{3}\tau_{3}+i\omega_{4}\tau_{4}-i\omega_{2}\tau_{2}-i\omega_{1}\tau_{1}}\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
+We also note that the Green’s functions contains frequencies of the driving field in their arguments.
+The expression can be simplified by plugging in the expressions of the exciton Green’s functions.
+It yields
+$\rho_{ff}(t)$ $\propto 2\mathrm{Re}\Big\{\int\frac{d\omega_{4}}{2\pi}\int\frac{d\omega_{3}}{2\pi}\int\frac{d\omega_{2}}{2\pi}\int\frac{d\omega_{1}}{2\pi}e^{+i\omega_{3}\tau_{3}+i\omega_{4}\tau_{4}-i\omega_{2}\tau_{2}-i\omega_{1}\tau_{1}}\langle E^{\dagger}(\omega_{4})E^{\dagger}(\omega_{3})E(\omega_{2})E(\omega_{1})\rangle$
 $\Big\{\frac{d_{fe^{\prime}}}{-\omega_{4}-\omega_{3}+\omega_{2}+\omega_{1}-z_{ff}}\frac{d_{e^{\prime}g}}{-\omega_{3}+\omega_{2}+\omega_{1}-z_{fe^{\prime}}}\frac{d_{fe}^{*}}{+\omega_{2}+\omega_{1}-z_{fg}}\frac{d_{eg}^{*}}{\omega_{1}-z_{eg}}$
@@
 $d_{e^{\prime}g}\left(\delta_{ee^{\prime}}\sum_{p}\chi^{\text{R}}_{e^{\prime\prime}p}D_{pp}^{-1}\frac{1}{-\omega_{3}+\omega_{1}-z_{p}}\chi^{\text{L}}_{pe^{\prime}}+(1-\delta_{ee^{\prime}})\delta_{ee^{\prime\prime}}\frac{1}{-\omega_{3}+\omega_{1}-z_{ee^{\prime}}}\right)\frac{d_{eg}^{*}}{\omega_{1}-z_{eg}}\Big\}\Big\}$ (26)
-A compact version of this expression, using the abbreviation $I(\omega_{j};z_{ab})=(\omega_{j}+z_{ab})^{-1}$ is presented in Eqn.
+A compact version of this expression, using the abbreviation $I(\omega_{j};z_{ab})=(\omega_{j}+z_{ab})^{-1}$ is presented in Eq.
 III.1.
-Proceeding via tedious yet straightforward frequency integrations yields the expression below
-$O_{ff}(t)$ $\propto 2\text{Re}\Big[\exp{(z_{ff}t)}\Big(d_{fe^{\prime}}d_{e^{\prime}g}d_{fe}d_{eg}^{*}\times\big\langle E^{\dagger}(z_{fe^{\prime}}-z_{ff})E^{\dagger}(z_{fg}-z_{fe^{\prime}})E(-z_{eg}+z_{fg})E(z_{eg})\big\rangle$
+In the final step, tedious yet straightforward frequency integrations yields the expression below
+$\rho_{ff}(t)$ $\propto 2\text{Re}\Big[\exp{(z_{ff}t)}\Big\{d_{fe^{\prime}}d_{e^{\prime}g}d_{fe}d_{eg}^{*}\times\big\langle E^{\dagger}(z_{fe^{\prime}}-z_{ff})E^{\dagger}(z_{fg}-z_{fe^{\prime}})E(-z_{eg}+z_{fg})E(z_{eg})\big\rangle$
 $+d_{fe^{\prime\prime}}d_{fe^{\prime\prime}}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\big\langle E^{\dagger}(+z_{fe^{\prime\prime}}-z_{ff})E(-z_{p}+z_{fe^{\prime\prime}})\sum_{p}\chi^{R}_{e^{\prime\prime}p}D_{pp}^{-1}E^{\dagger}(z_{eg}-z_{p})\chi^{L}_{pe}E(z_{eg})\big\rangle$
@@
 $+d_{fe^{\prime\prime}}d_{fe^{\prime\prime}}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\big\langle E^{\dagger}(-z_{e^{\prime\prime}f}+z_{ff})E(z_{p}-z_{e^{\prime\prime}f})\sum_{p}\chi^{R}_{e^{\prime\prime}p}D_{pp}^{-1}E^{\dagger}(z_{ee^{\prime}}-z_{p})\chi^{L}_{pe}E(z_{eg})\big\rangle$
-$+d_{fe^{\prime\prime}}d_{fe^{\prime\prime}}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\big\langle E^{\dagger}(-z_{ef}+z_{ff})E(+z_{ee^{\prime}}-z_{ef})E^{\dagger}(z_{eg}-z_{ee^{\prime}})E(z_{eg})\big\rangle\Big)\Big]$ (27)
-which is amenable to numerical simulation.
-Appendix C Entangled photon correlation function
-Here we present the numerical simulation of the square of the joint spectral amplitude function, for typical parameter regime used in the main simulation.
+$+d_{fe^{\prime\prime}}d_{fe^{\prime\prime}}^{*}d_{e^{\prime}g}d_{eg}^{*}\times\big\langle E^{\dagger}(-z_{ef}+z_{ff})E(+z_{ee^{\prime}}-z_{ef})E^{\dagger}(z_{eg}-z_{ee^{\prime}})E(z_{eg})\big\rangle\Big\}\Big]$ (27)
+This expression is valid for any general correlated photonic source and can be used in numerical simulation.
+Appendix D Entangled photon correlation function
+Here we examine the time-frequency properties entangled photon pairs.
+We plot the square of the joint spectral amplitude function, for a parameter regime comparable to that used in the simulation in Section III.
 Figure 10: The squared joint spectral amplitudes of the entangled photon pairs are displayed, illustrating variations in the entanglement time parameter ($\tilde{T}_{\text{ent}}$) and the temporal width of the SPDC pump ($\tau_{0}$).
 They have noticeable implications for the nature of the frequency (anti-)correlations between the signal and idler pair, a feature that can be utilized for spectroscopic applications.
-Appendix D Temporal spectral filtering functions
-We use the following temporal ${F}_{\bar{t}}(t)$, and spectral $F_{\bar{\omega}}(\omega)$ filtering functions, expressed as
-$F_{\bar{t}}(t^{\prime})$ $=\theta(t^{\prime}-\bar{t})\exp({-\sigma_{\bar{t}}(t^{\prime}-\bar{t})})$ (28)
-$F_{\bar{\omega}}(\omega^{\prime})$ $=i(\omega^{\prime}-\bar{\omega}+i\sigma_{\omega})^{-1}$ (29)
-These functions have four tunable parameters: the centering variables, $\bar{t}$, $\bar{\omega}$, and the width $\sigma_{\bar{t}}$, $\sigma_{\bar{\omega}}$.
-The presented simulations in Fig.
-7 illustrates their role in signal classification.
-Appendix E Note on the observable: photon coincidence signal
-Here, we present a brief outline of derivation of the bare photon coincidence counting observable (i.e., without the filtering stage), following the original theory developed in [13, 11].
-The expression is already presented in Eq.
-III, and appears as
+Appendix E Note on the observable: two-photon counting signal
+Here, we present a brief derivation of the photon coincidence counting signal. We follow the original steps outlined in [17, 15]. The final expression is already presented in Eq. III, and appears as
+The bare two-photon counting signal is expressed as
 $\langle N_{s}(t_{1}^{\prime},\tau_{1})N_{i}(t_{2}^{\prime},\tau_{2}^{\prime})\rangle=\int_{-\infty}^{t_{1}^{\prime}}dt_{1}\int_{-\infty}^{t_{1}^{\prime}+\tau_{1}}dt_{3}$
@@
 $\times E_{iL}(t_{1}^{\prime})E_{sL}(t_{2}^{\prime})E_{iL}^{\dagger}(t_{1})E_{sL}^{\dagger}(t_{2})\rangle.$ (28)
-Here, the averaging of the field correlation term is carried out over the detector modes, and the auxiliary photon modes, where as the averaging of the exciton correlation term is carried out over two-exciton state following transport.
-In the presented theory, the leading order contributions to the signal arise from expansion of the time dependent density operator to the fourth order in emitted field.
-Assuming photon emission into vacuum modes, and following the standard algebraic manipulations one obtains
+The field correlation term is averaged over the detector modes in vacuum.
+The exciton correlation function is averaged over the the two-exciton states that corresponds to the population terms of the density operator following the transport.
+We assumed that the leading order contributions to the photon counting signal arise from expansion of the density operator to the fourth order in emitted photon field.
+The emitted field modes are expanded into plane wave mode and following the standard algebraic manipulations we obtain
 $\langle N_{s}(t_{1}^{\prime},\tau_{1})N_{i}(t_{2}^{\prime},\tau_{2}^{\prime})\rangle=\mathcal{D}^{2}(\omega_{1})\mathcal{D}^{2}(\omega_{2})$
 $\times\langle V_{R}^{\dagger}(t_{2}^{\prime}+\tau_{2})V_{R}^{\dagger}(t_{1}^{\prime}+\tau_{1})V_{L}(t_{1}^{\prime})V_{L}(t_{2}^{\prime})\rangle.$ (29)
-This expression represents the exciton correlation function weighted by the density of detector modes.
-The correlation function can be further expanded in terms of individual pathways Fig.
-2, yielding the expressions in Eq.
-III.3.
+It highlights the fact that the four-point exciton correlation function weighted by the density of detector modes evaluated at emitted frequencies.
+The exciton correlation function can be expanded further to yield the Liouville space pathways depicted in the diagram in Fig.
+3.
+The expansion also yields the expressions in  III.3.